### v1

# HyGen: Efficient LLM Serving via Elastic Online-Offline Request Co-location

Abstract
Recent advancements in large language models (LLMs) have facilitated a wide range of applications with distinct quality-of-experience requirements, from latency-sensitive online tasks, such as interactive chatbots, to throughput-focused offline tasks like document summarization. While deploying dedicated machines for these services ensures high-quality performance, it often results in resource underutilization.
This paper introduces HyGen, an interference-aware LLM serving system that enables efficient co-location of online and offline workloads while preserving latency requirements. HyGen incorporates two key innovations: (1) performance control mechanisms, including a latency predictor for batch execution time estimation and an SLO-aware profiler to quantify interference, and (2) SLO-aware offline scheduling policies that maximize throughput and prevent starvation, without compromising online serving latency. Our evaluation on production workloads shows that HyGen achieves up to 5.84$\times$ higher throughput compared to existing advances while maintaining comparable latency. The code will be released soon.
1 Introduction
Large language models (LLMs) have emerged as transformative tools across diverse domains, powering applications from interactive chatbots to document analysis systems Grand View Research (2023). As LLM services continue to spread, their serving deployment must handle both latency-critical online requests (e.g., chatbot interactions OpenAI (2022); Liu et al. (2024a)) and throughput-oriented offline tasks (e.g., document summarization Grand View Research (2023)). Online serving demands low and stable response times, measured by Time to First Token (TTFT) and Time Between Tokens (TBT), to ensure smooth user interactions. In contrast, offline tasks prioritize high throughput and resource utilization, often processing large batches of requests with relaxed latency constraints. The disparity in these requirements has led most production deployments to segregate online and offline serving onto separate GPU clusters Sheng et al. (2023); Zhao et al. (2024); Patel et al. (2024); Stojkovic et al. (2024).
However, real-world LLM workloads demonstrate significant temporal variations in request loads. Our analysis of production traces shows that in addition to the popular diurnal request arrival patterns, online request rates can vary by up to 3$\times$ within minutes (Section 3). To maintain acceptable latency under such bursty loads, service providers must provision GPU resources for peak demand, leading to substantial underutilization during off-peak hours.
This resource overprovisioning suggests an opportunity to improve resource efficiency via hybrid serving—co-locating online and offline workloads on the same GPU cluster. By opportunistically co-locating offline tasks with online tasks during periods of low online load, the system could maintain high GPU utilization while preserving latency guarantees for online requests. However, realizing this opportunity requires addressing several fundamental challenges.
First, LLM services exhibit diverse latency requirements across applications Liu et al. (2024a). Interactive chatbots require consistent response times with both low initial latency (TTFT) and smooth token generation (TBT), while batch processing tasks prioritize throughput over latency. These requirements often manifest in different statistical metrics—from strict P99 latency bounds to mean performance targets—making it difficult to establish unified resource-sharing policies.
Second, LLM workloads are inherently unpredictable in both their arrival patterns and resource demands. Online requests arrive in unexpected bursts, while offline tasks appear in irregular batch intervals. This variability is further complicated by uncertainty in resource demands—input sequences vary widely in length, and the number of output tokens can hardly be predicted until generation completes.
Third, co-locating online and offline workloads introduces interference, which degrades the performance of both. For instance, offline requests can delay time-sensitive online requests, while using large batch sizes to boost offline request throughput can negatively impact the latency of online requests. Effectively managing this interference while ensuring performance guarantees demands meticulous orchestration of resource sharing.
This paper presents HyGen, an interference-aware LLM serving system that elastically co-locates online and offline workloads. HyGen introduces several key techniques: (1) a latency predictor that accurately estimates the execution time of heterogeneous request batches, (2) an interference-aware profiler that quantifies the performance impact of co-location, and (3) an adaptive scheduler that maximizes offline throughput while maintaining strict latency guarantees for online serving.
Our evaluation on production workloads shows that HyGen guarantees latency requirements in co-locating requests and achieves up to 5.84$\times$ higher throughput compared to existing advances while maintaining comparable latency.
To summarize, this paper makes the following contributions:
  • We investigate the online request patterns in production LLM serving trace, revealing key challenges and opportunities in co-locating online and offline workloads.
  • We propose HyGen, a new serving system that enables efficient resource sharing, improving throughput with guaranteed performance isolation.
  • To improve the performance isolation and throughput of HyGen, we design performance control mechanisms and SLO-aware offline scheduling policies.
  • We evaluate HyGen using production LLM serving traces to demonstrate HyGen’s effectiveness.
2 Related Work
LLM Online Serving.
Recent works have improved GPU computation through optimized kernels Dao et al. (2022); Ye et al. (2025), memory efficiency through efficient KV-cache management techniques Lin et al. (2024a); Liu et al. (2024b); Qin et al. (2024), distributed inference through various model parallelism strategies Shoeybi et al. (2019); Pope et al. (2023); Liu et al. (2023), and model compression through quantization approaches Xiao et al. (2023); Lin et al. (2024b); Dettmers et al. (2022). Orca Yu et al. (2022) introduces iteration-level scheduling to eliminate padding overhead in variable-length sequence batching, and enables early exits for completed requests, thereby reducing batching latency. vLLM Kwon et al. (2023) develops PagedAttention for efficient memory management, while Sarathi-Serve proposes chunked prefill to minimize inter-request interference within batches. However, few existing systems explicitly address the challenge of co-locating latency-critical online requests with offline workloads to maximize resource utilization—a key focus of our work.
Throughput-Oriented LLM Inference.
As offline inference gains increasing traction, Huggingface Accelerate Hugging Face (2025), DeepSpeed ZeRO-Inference Aminabadi et al. (2022) and FlexGen Sheng et al. (2023) enable efficient LLM serving on commodity GPUs by offloading strategies. FlexGen aggregates memory from GPU, CPU and disk to maximize serving throughput. For KV cache optimization, InfiniGen Lee et al. (2024) proposes dynamic cache management through selective prefetching based on token importance prediction. S3 Jin et al. (2023) tackles the KV cache problem through output sequence length prediction, allowing more efficient memory allocation by avoiding over-allocation of the maximum sequence length. NanoFlow Zhu et al. (2024) focuses on maximizing hardware utilization through intra-device parallelism, introducing nano-batching and execution unit scheduling to overlap compute-, memory-, and network-intensive operations.
Co-locating Workloads.
Co-locating latency-critical applications with batch applications is a widely adopted approach to improve resource utilization in datacenters, with the popular resource partitioning techniques Chen et al. (2019b); Zhu & Erez (2016); Lo et al. (2015); Wu & Martonosi (2008); Chen et al. (2019a); Kasture et al. (2015); Sanchez & Kozyrakis (2011); Kasture & Sanchez (2014); Zhang et al. (2019); Iorgulescu et al. (2018); Patel & Tiwari (2020); Chen et al. (2023). Prior work has also explored multiplexing to serve multiple DNNs Dhakal et al. (2020); Tan et al. (2021); Zhang et al. (2023b), LLMs Duan et al. (2024), or LoRAs Chen et al. (2024); Sheng et al. (2024a); Wu et al. (2024). However, prior efforts have largely overlooked the opportunity to co-locate online and offline inference requests within the same LLM.
3 Background and Motivation
This section first introduces the core concepts of LLM serving systems (Section 3.1), then illustrates how these characteristics motivate our system designs (Section 3.2).
3.1 LLM Serving
LLM generation consists of two sequential stages: (i) Prefilling Stage: It processes the entire input prompt in parallel to generate the first output token, computing and caching intermediate model representations (i.e., the KV cache) of the input sequence to avoid regeneration in subsequent tokens, and then the (ii) Decoding Stage: the model generates tokens one at a time in an auto-regressive manner. Each iteration processes only the most recent token while utilizing the KV cache of previous tokens, reducing computational overhead. This stage iterates until meeting a termination condition, such as generating an end-of-sequence token or reaching the maximum length.
Model executions in the prefilling stage are compute-intensive due to the high degree of parallelism in processing the entire input sequence simultaneously. Yet in the decoding phase, the same operations become memory-intensive due to the reliance on the KV cache Agrawal et al. (2024), and the total memory access cost grows linearly with the sequence length as each new token’s KV cache is appended.
Serving Workloads.
LLM serving deployment can be categorized into two primary scenarios. The first scenario is online serving, which targets real-time user interactions, such as chatbots, code assistants, and interactive applications OpenAI (2022); Microsoft (2023). This interactive nature often requires short Time to First Token (TTFT)—the latency from receiving a query to generating the first token—as well as the short Time Between Tokens (TBT), i.e., the inter-token generation latency in the decode stage.
The second counterpart is offline serving, which prioritizes throughput over latency, measured in requests per second (RPS) or tokens per second (TPS) for prefill/decode operations. Key applications include model benchmarking Hendrycks et al. (2021b); Meta-Team (2024), document processing Chen et al. (2021); Zhang et al. (2023a), data cleaning Narayan et al. (2022); Zheng et al. (2023), and synthetic data generation Allal et al. (2024).
3.2 Motivation and Challenges
(a) 1 hour
(b) 2 minutes
Figure 1: Request rate variation in Microsoft Azure’s LLM service over one-hour and two-minute periods, showing significant temporal fluctuations between peak and off-peak periods.
LLM serving systems face a critical resource utilization challenge due to the highly variable nature of their workloads. Figure 1 reports our analysis of Microsoft Azure’s LLM service Patel et al. (2024), showing that request rates can fluctuate dramatically—varying up to 3× within minutes while following broader diurnal patterns. This variability creates an inherent tension in resource provisioning: GPU clusters must be sized to handle peak loads, inevitably leading to resource underutilization during off-peak periods.
This load variability suggests an opportunity to improve GPU utilization by co-locating online inference requests with offline batch workloads. During periods of low online traffic, the system could opportunistically schedule offline tasks to harvest idle GPU resources. However, co-locating these workloads introduces several fundamental challenges:
  1. 1.
    Diverse Latency Requirements: Applications and even requests of an application have distinct latency requirements. For example, users of different ages have different reading speeds, thus TBT requirements Liu et al. (2024a). These requirements are often expressed through multiple statistical metrics - some services prioritize tail latency guarantees (e.g., P99 TTFT $\leq$ 1000ms), others focus on average performance (e.g., mean TBT $\leq$ 50ms), and many need to satisfy both simultaneously. How to provide latency guarantees for diverse requests while sharing resources?
  2. 2.
    Massive Uncertainties: LLM serving faces temporal uncertainty—online requests arrive in unpredictable bursts with varying urgency levels—as well as resource demand uncertainty due to unpredictable output lengths. A seemingly simple query might trigger an elaborate response requiring thousands of tokens. How to perform efficient scheduling in the wild?
  3. 3.
    Request Interference: Co-locating online and offline workloads introduces interference that degrades system performance. Large batches of offline requests arriving simultaneously can cause severe head-of-line blocking, delaying the processing of time-sensitive online requests. Worse, batching requests of long inputs (e.g., a long document summarization request with 2,048 tokens) with short interactive queries (e.g., 128 tokens) can result in a dramatic increase in attention computation overhead, elongating the latency for all requests in the batch by an order of magnitude Agrawal et al. (2024). How to account for interference in co-locating requests?
4 The HyGen Design
4.1 Overview
Figure 2: HyGen overview: Online requests and offline requests sent by users are asynchronously selected and processed.
In this paper, we present HyGen, a novel LLM serving system that enables efficient co-location of online and offline requests while maintaining latency guarantees. Unlike conventional LLM serving systems that segregate online and offline workloads onto separate GPU clusters, HyGen coordinates both types of requests through a two-phase scheduling approach, maximizing resource utilization during off-peak periods while preserving performance isolation.
Algorithm 1 HyGen SLO-aware scheduler
1:  function SLO_AWARE_SCHEDULE
2:     Input: running requests $R$, request queue $Q$,
3:      latency budget $t$, remaining chunk size $c$,
4:      memory budget $m$
5:     Output: batched requests $B$
6:     $B\leftarrow\{\}$
7:     for $r\in R$.decoding do
8:        $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
9:        if $t_{req}\leq t$ then
10:           $t\leftarrow t-t_{req}$
11:           $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
12:        end if
13:     end for
14:     for $r\in R$.prefilling $\cup Q$ do
15:        TRY_SCHEDULE:
16:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
17:        if $l>0$ then
18:           $t\leftarrow t-t_{req}$
19:           $c\leftarrow c-l$
20:           $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
21:           $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
22:        else
23:           if PHASE ==absent$=$ = ONLINE and $R\neq B$ then
24:              PERFORM_PREEMPTION($R$, $m$)
25:              goto TRY_SCHEDULE
26:           else
27:              break
28:           end if
29:        end if
30:     end for
31:     return $B$
32:  end function
Figure 2 illustrates the system architecture of HyGen. To handle diverse latency requirements and unpredictable workloads, HyGen maintains separate queues for online and offline requests, each managed by a dedicated request selector that tracks active requests and their states. This separation allows HyGen to support scheduling policies for different request types without the complexity of managing mixed priorities in a single queue.
The scheduling process in HyGen operates in two consecutive phases to carefully manage interference between online and offline requests. In the online phase, the scheduler processes the online request queue to construct an initial batch based on latency SLOs. This phase follows traditional continuous batching principles and supports various scheduling policies such as First-Come-First-Serve (FCFS) Kwon et al. (2023) and Virtual Token Counter (VTC) Sheng et al. (2024b). The scheduler prioritizes running requests before processing the waiting queue, continuing until reaching memory, batch size, or chunk size limits. A detailed workflow of this process is displayed in Algorithm 1. During a default two-phase scheduling, this function is called two times separately for online and offline requests.
During online batch establishment, when the current online batch reaches the memory limit and more online requests remain to be scheduled, HyGen performs a priority-based preemption to maximize SLOs. To handle dynamic request arrivals while maintaining service guarantees, we implement an online-first preemption mechanism, where online requests are prioritized over all offline requests to minimize interference. Offline requests are prioritized based on their original priority in the waiting queue and their memory consumption. Preempted offline requests are removed from the current running list and reinserted into the offline queue, where they maintain their execution state to enable efficient rescheduling in subsequent batches.
After establishing the online batch, the scheduler enters the offline phase where it opportunistically fills the remaining capacity with offline requests. By utilizing a latency predictor to estimate batch execution time and determine remaining latency budget, instead of relying on static chunk size or batch size limits, we achieve both higher accuracy in interference prediction and strong transferability across different hardware configurations and workload patterns. This enables higher throughput while maintaining SLO guarantees by eliminating the need for conservative scheduling. The selection of offline requests then considers this available batch execution latency budget derived from user-specified SLOs, opportunities for efficient execution through prefix sharing, and current batch composition and resource constraints. Within this phase, running requests are prioritized over waiting ones to maintain execution efficiency.
HyGen addresses these complex scheduling decisions through two key components detailed in the following sections. Section 4.2 describes our performance control mechanisms, including the latency predictor and SLO-aware profiler that enable accurate resource allocation decisions. Section 4.3 presents our SLO-aware offline scheduling policies that maximize throughput while maintaining performance isolation through careful workload co-location. A detailed algorithm for the entire HyGen scheduling workflow can be found in Appendix A.1.
4.2 Performance Control Mechanisms
Efficient co-location of online and offline requests requires precise control over execution timing and resource allocation. The key challenge lies in accurately estimating the performance impact of scheduling decisions while ensuring end-to-end SLO compliance. This section presents two core mechanisms in HyGen that address this challenge: a latency predictor for estimating batch execution times and an SLO-aware profiler for translating these estimates into scheduling decisions.
Latency Predictor.
The design of our latency predictor is guided by three key requirements. First, it must provide fast inference to support real-time scheduling decisions when evaluating different batch compositions. Second, it needs to be robust across varying workload patterns conditions to maintain reliable performance. Third, it should be adaptable to different hardware configurations to capture their unique performance characteristics.
The execution time of an LLM serving batch is primarily influenced by two distinct processing stages with different computational patterns. The prefill stage exhibits quadratic complexity due to attention computations, with latency growing quadratically with input sequence length. This stage’s total load depends on both the number of requests and their individual sequence lengths. In contrast, the decode stage shows linear scaling, with computational requirements growing proportionally with the number of tokens.
To capture these characteristics, we model batch execution time as:
$T_{batch}=f(S_{p},S_{d},S_{p}^{2},S_{d}^{2},N_{p},N_{d})$ (1)
where $S_{p}$ and $S_{d}$ represent the total number of prefill and decode tokens in the batch, respectively. The quadratic terms ($S_{p}^{2}$ and $S_{d}^{2}$) account for non-linear scaling effects, particularly in the prefill phase. $N_{p}$ and $N_{d}$ represent the number of prefill and decode requests in the batch, respectively.
We implement this model using linear regression, chosen for its combination of efficiency and effectiveness. The linear model enables rapid evaluation of different batch compositions during scheduling, while its simple feature set ensures stable predictions across varying conditions. The training process involves three steps:
  1. 1.
    Training Data Collection: We systematically profile target hardware across diverse batch compositions, varying the number of requests in different phases, sequence length distributions, and total batch sizes.
  2. 2.
    Feature Engineering: For each batch, we compute six key features ($S_{p}$, $S_{d}$, $S_{p}^{2}$, $S_{d}^{2}$, $N_{p}$, $N_{d}$) based on token counts and batch sizes, capturing both linear and quadratic scaling.
  3. 3.
    Model Training: Linear regression coefficients are optimized to minimize prediction error across the training set, providing sufficient accuracy for scheduling decisions while maintaining fast inference times.
SLO-Aware Profiling.
While accurate latency prediction is essential, translating these predictions into scheduling decisions requires careful consideration of end-to-end performance requirements. Our SLO-aware profiler addresses this by determining appropriate latency budgets that ensure SLO compliance under various operating conditions.
The profiler first analyzes specific workload-SLO combinations to establish viable latency budget ranges. Then, given that SLO has a positive correlation with batch latency, we perform the binary search on the batch latency budget to get an upper limit that meets the SLO. This profiling enables three key capabilities. First, it determines appropriate latency thresholds that maintain SLO compliance for different workload types. Second, it provides flexible adaptation by adjusting budgets based on changing workload characteristics and performance requirements. Third, it establishes a robust foundation for hybrid scheduling by accounting for both online and offline workload patterns.
Together, these mechanisms enable HyGen to make precise scheduling decisions that maximize resource utilization while maintaining strict performance guarantees. The latency predictor provides the quantitative basis for evaluating scheduling options, while the SLO-aware profiler ensures these decisions translate into consistent end-to-end performance.
4.3 SLO-Aware Offline Scheduling Policies
Offline scheduling in HyGen differs from online scheduling due to the variability of resource availability. Resources allocated to offline requests depend on the residual capacity after scheduling online requests, which fluctuates with the intensity of the online workload. Consequently, offline scheduling policies must dynamically adapt to available resources and optimize their specific objectives under such constraints. Common objectives include maximizing throughput, supporting multi-priority workloads, and, optionally, ensuring fairness Sheng et al. (2024b). The prioritization of these objectives varies depending on the chosen policy.
Prefix sharing Gim et al. (2024); Gao et al. (2024); Zheng et al. (2024a) has emerged as a widely-adopted technique to significantly improve throughput and reduce latency in LLM serving systems. The benefits are particularly pronounced in offline LLM serving, where requests can be strategically ordered to maximize prefix cache reuse Zhao et al. (2024); Zheng et al. (2024b). This optimization is especially effective for offline workloads since they have relaxed latency constraints, allowing the system to accumulate and reorder requests to maximize sharing opportunities. To showcase the extensibility and flexibility of HyGen’s design, we next introduce the SLO-aware Prefix Sharing Maximization (PSM) strategy, a throughput-optimized offline scheduling policy that leverages prefix reuse to improve resource efficiency. Recognizing the potential starvation issue inherent in naive PSM, we extend it with a fairness-enhanced variant. This improved PSM algorithm incorporates mechanisms such as freshness-aware scheduling and a unified utility function to balance throughput and fairness, mitigating starvation without compromising efficiency.
The modular architecture of HyGen enables seamless integration of advanced offline scheduling policies. In addition to PSM, the system supports other strategies, such as a default First-Come-First-Serve (FCFS) policy and a Deadline-First policy for basic scheduling needs. This flexibility allows HyGen to cater to a wide range of objectives, including multi-priority support and fairness, making it a robust and adaptable solution for hybrid serving scenarios with dynamic and heterogeneous workloads.
Prefix Sharing Maximization Strategy.
To fully leverage prefix sharing among offline requests, HyGen employs a Prefix Sharing Maximization (PSM) strategy designed for SLO-aware offline scheduling. This strategy organizes offline requests into a prefix tree and performs Depth-First Search (DFS) to generate a sorted waiting list. HyGen then performs SLO-aware scheduling on offline requests using this sorted list in a similar manner as introduced in Section 4.1, while maintaining the original DFS order to maximize prefix sharing. This enables reducing redundant computation and improving throughput.
Despite its efficiency, the naive PSM strategy can lead to starvation for requests with minimal prefix-sharing potential. This issue can be mitigated by combining requests selected by PSM with those chosen based on freshness. The trade-off between these two selection methods can be managed using a utility ratio, ensuring a balance between efficiency and fairness. During scheduling, offline waiting requests are stored in both a prefix tree for maximum prefix sharing, and a self-balanced binary search tree (BST) sorted by freshness. When a new offline request can be scheduled from the waiting list, an offline request is selected either from the head of the queue in the DFS order of the prefix tree, or from the most stale request in the self-balanced BST, based on the utility value. The other data structure then syncs up by deleting the selected request from their lists, namely by prefix and index lookup. The formalized algorithm is detailed in Appendix A.2.
While starvation due to the strict prioritization of online requests over offline ones is unavoidable to ensure performance isolation, the enhanced PSM algorithm effectively mitigates starvation caused by prefix-sharing optimizations. These enhancements make the PSM strategy more practical for real-world deployments, retaining its efficiency while improving fairness and adaptability.
5 Performance Evaluation
Figure 3: HyGen respects latency requirements in co-locating requests.
Figure 4: HyGen improves serving throughput under the same SLOs.
5.1 Evaluation Setup
Implementation.
We implement HyGen on top of the open-source implementation of vLLM Kwon et al. (2023); vLLM Project (2023) and Sarathi-Serve Agrawal et al. (2024); Sarathi-Serve Project (2024), with 1,300 lines of additional code.
Testbeds.
We evaluate HyGen on two server configurations. The first server is equipped with 4 NVIDIA A100 GPUs (40GB VRAM each), while the second server contains 4 NVIDIA A40 GPUs (48GB VRAM each). Both servers are identically configured with an AMD EPYC Milan CPU (1 socket, 64 cores, SMT disabled), 256GB DDR4 RAM, and a 1.5TB NVMe SSD (PCIe 4.0).
Models.
We used Llama2-7B Touvron et al. (2023) and Qwen-14B Bai et al. (2023) models to test the performance of HyGen on varying model structures and sizes. For the two models, we used A100 and A40 GPUs, respectively.
Workloads.
For the online workload, we used the conversation trace from Azure LLM inference trace 2023 Patel et al. (2024), a one-hour trace taken from the LLM services in Azure with real-world requests and timestamps. The request rate variation is shown in Figure 1. We randomly sampled the original trace to get the desired overall QPS that suits our hardware serving capacity. Specifically, for a time duration of $T$ seconds, we would sample $T\times Q$ requests from this time window to suit a desired QPS $Q$. For the offline workload, we used arXiv summarization Cohan et al. (2018), a dataset for summarization of long documents.
Metrics.
For interference evaluation, we focus on both Time to First Token (TTFT) and Time Between Tokens (TBT) and their respective mean and 99th percentile (P99) values. For throughput evaluation, we evaluated our system on the maximum throughput of offline requests while meeting interference SLOs.
Baselines.
For pure online inference, we use Sarathi Agrawal et al. (2024) as our baseline. For hybrid serving, we compare HyGen with two baselines for interference and throughput evaluation, respectively: (1) Sarathi++: We implement our online-first two-phase scheduling policy on Sarathi to support hybrid serving, including the request management and preemption handling policies introduced in Section 4. (2) HyGen*: To evaluate the throughput benefit of our HyGen design, we further improve Sarathi++ to be an SLO-aware serving system, HyGen*. Besides inheriting the serving policies from Sarathi++, HyGen* serves offline requests at a specific offline QPS to control overall interference. The offline QPS is profiled using a similar design with the HyGen profiler to guarantee bounded SLO interference.
5.2 End-to-end Performance
HyGen respects latency requirements in co-locating requests. We tested HyGen on four SLO metrics, namely mean TBT, P99 TBT, mean TTFT, and mean TBT. For each metric, we set varying interference tolerance ratios as the SLO. Figure 3 shows that HyGen is able to control the interference and guarantees to meet specific SLOs across our settings. Compared with Sarathi++, an SLO-unaware system that yields the same result for all metrics and tolerance ratios, HyGen showcases efficient SLO-aware latency control.
Figure 5: Prefix sharing maximum scheduler achieves higher serving capacity.
Figure 6: SLO-aware profiler bridges the gap between batch latency and overall QoE.
Figure 7: HyGen latency predictor achieves high accuracy for batch latency prediction.
Figure 8: HyGen meets stringent SLO under various QPS settings
Figure 9: Hygen dynamically controls offline throughput according to online workload and latency budget.
(a) SLOs
(b) Throughput
Figure 10: Hygen meets SLOs and achieves higher throughput for Yi-34B model using TP=2, PP=2.
(a) SLOs
(b) Throughput
Figure 11: Hygen meets SLOs and achieves higher throughput for Mooncake trace and Mistral-7B model.
(a) SLOs
(b) Throughput
Figure 12: Hygen meets SLOs and achieves higher throughput for CNN/DM offline trace.
Figure 13: HyGen is able to meet multiple SLOs.
Figure 14: HyGen design is robust to predictor accuracy.
Figure 15: Hygen adjusts offline throughput based on online QPS.
HyGen improves serving throughput. Figure 4 showcases the offline throughput of HyGen for various metrics and tolerance ratios. Under the same SLO, HyGen consistently achieves higher throughput compared to HyGen*, yielding up to 5.37×5.37$\times$5.37 × to 5.84×5.84$\times$5.84 × throughput gain. This verifies that our fine-grained latency predictor and SLO-aware profiler achieve higher offline serving efficiency compared to their simplified counterparts.
5.3 Performance Breakdown
Accuracy of latency predictor.
We evaluate the accuracy of our latency predictor on Llama2-7B and Qwen-14B using Azure LLM Inference trace mixed with arXiv summarization dataset. Figure 7 shows the CDF of absolute error and absolute percentage error of the latency predictor. HyGen latency predictor is able to achieve a mean absolute percentage error of only 1.78% and 1.07%, verifying the high accuracy of our predictor design.
Impact of prefix sharing.
To test HyGen’s compatibility with prefix sharing, we did a simulation experiment using Azure LLM Inference as the online trace and MMLU Hendrycks et al. (2021a; b) as the offline dataset, running on Llama2-7B model. In our simulation, the shared prompt prefix length between two consecutively scheduled offline requests is deducted from the original prompt length of the latter one to simulate prefix sharing. We tested the offline throughput of HyGen with and without our prefix-aware offline schedule policy. Figure 6 shows that HyGen yields up to 4×4$\times$4 × offline throughput gain with its prefix-aware scheduling policy.
Impact of SLO-aware profiler.
To demonstrate the effect of the SLO-aware profiler in HyGen, we conducted experiments where we simply set the desired mean TBT SLO as the batch latency budget. Figure 6 shows the performance gap between individual batch latencies and overall mean TBT, and that our SLO-aware profiler is able to bridge this gap and provide hybrid serving with controlled SLO.
Breakdown by time.
Figure 9 shows a temporal breakdown of the throughput performance of HyGen during one serving process. At runtime, HyGen shows the ability to dynamically adjust offline throughput according to current online workload and overall latency budget. HyGen batches offline requests more aggressively during troughs of online QPS and lowers offline throughput during online bursts, harnessing compute resources in an adaptive manner.
5.4 Ablation Studies
Impact of model parallelisms.
To evaluate HyGen’s effectiveness in distributed inference scenarios, we deployed the Yi-34B model Young et al. (2024) using a hybrid parallelization strategy: tensor-parallelism (TP) Shoeybi et al. (2019) and pipeline-parallelism (PP) Athlur et al. (2022); Narayanan et al. (2019); Huang et al. (2019) with degree 2 for each dimension. The online and offline workloads are Azure LLM Inference trace and arXiv summarization, respectively. Figure 10(a) and Figure 10 show that HyGen maintains its ability to meet SLOs and achieves higher throughput (up to 1.89×1.89$\times$1.89 ×) than the baseline.
Impact of SLO requirements.
To further evaluate HyGen’s ability to meet stringent SLOs under various settings, we conducted experiments on the four aforementioned metrics under varied QPS settings, each with a 5% interference tolerance ratio. Figure 8 shows that HyGen is able to meet stringent SLOs for all metrics under various QPS settings.
We further demonstrate HyGen’s ability to meet multiple SLOs at the same time. Here, we tested HyGen with a fixed P99 TTFT interference ratio (8%) and mean TBT interference ratios ranging from 10% to 50%. Figure 13 shows that at a lower mean TBT tolerance, the performance of HyGen is bounded by mean TBT SLOs; however, after reaching the fixed P99 TTFT SLO, mean TBT stops increasing in order to keep P99 TTFT under control, demonstrating the flexibility of HyGen in meeting multiple SLOs.
Impact of models and datasets.
We further evaluate HyGen on Mistral-7B and MooncakeQin et al. (2024), a trace containing request lengths and timestamps taken from real-world servers, as the online trace. The offline dataset used in the first experiment is arXiv summarization. For the second experiment, we replaced the offline dataset with CNN/DailyMail summarization dataset See et al. (2017); Hermann et al. (2015) and used Llama2-7B and Azure LLM Inference trace as the model and online dataset, respectively. Figure 11 and Figure 12 show HyGen achieves superior performance than its counterparts under these two settings.
Impact of predictor accuracy.
To further evaluate the robustness of HyGen, we tested HyGen using several pre-trained LR latency predictors taken from other settings and workloads and tested them on Azure LLM Inference trace, resulting in varied prediction accuracy. Figure 15 shows how predictor accuracy (measured in mean absolute percentage error) affects offline throughput under the same P99 TBT SLO. We notice that HyGen remains robust across different accuracy settings. Additionally, our LR-based latency predictor is lightweight for training, with only $\sim$15ms training time for over 80,000 samples on our testing CPU, guaranteeing transferability and negligible overhead.
Impact of online arrival rates.
Under high online request arrival rates, online requests may approach system serving capacity, leaving little headroom for offline request co-location. Figure 15 shows how online QPS affects offline throughput under a 5% P99 TBT SLO tolerance ratio. As online load increases, HyGen is able to adjust offline throughput according to the residual serving capacity of the system, while maintaining higher throughputs.
6 Conclusion
This paper introduces HyGen, an LLM serving system that enables efficient co-location of online and offline workloads. We employ a control mechanism to predict and manage interference impacts, while the scheduling framework supports multiple optimization objectives for offline serving. Evaluation on production workloads demonstrates that HyGen achieves up to 5.84$\times$ higher throughput while maintaining strict latency SLOs.
Appendix
Appendix A Algorithms
A.1 HyGen Two-phase Scheduling Algorithm
This section gives a detailed explanation of the two-phase scheduling algorithm introduced in Section 4. To reduce the scheduling overhead of HyGen, we employ a message queue for asynchronous communication between the main process and the offline scheduler. After each scheduling step, the main process sends the metadata of batched requests to the message queue. The offline scheduler first calculates the expected status of each request based on scheduling decisions from the previous batch, and then runs a scheduling simulation to generate offline request scheduling decisions using our latency predictor and the profiled latency budget. The offline scheduling decisions are then sent back to the main process using the message queue and used for the coming scheduling step. To support pipeline parallelization, a scheduling history archive of $K$ steps is kept by the offline scheduler for pipeline parallelization size $K$, in order to have a holistic view of every request running in each pipeline stage at the time.
Algorithm 2 HyGen two-phase scheduler
1:  global $Q_{on}$ (online request queue), $Q_{off}$ (offline request queue)
2:  global $R_{on}$ (online request running list), $R_{off}$ (offline request running list)
3:  global $Q_{send}$, $Q_{recv}$ (message queue)
4:  global $L$ (latency budget), $M$ (memory budget), $C$ (chunk size), $M_{off}$ (offline memory)
5:  function ASYNC_SCHEDULER
6:     while True do
7:        scheduled requests $S\leftarrow Q_{send}$.get(block=True)
8:        UPDATE_REQUEST_STATUS($S$, $R_{on}$, $R_{off}$)
9:        batched requests $B\leftarrow$}
10:        latency budget $t\leftarrow L$
11:        memory budget $m\leftarrow$ GET_FREE_MEMORY() + $M_{off}$
12:        chunk size $c\leftarrow C$
13:        $B\leftarrow B\cup$ SLO_AWARE_SCHEDULE($R_{on}$, $Q_{on}$, $t$, $c$, $m$)
14:        if $m<M_{off}$ then
15:           PREEMPT_OFFLINE($R_{off}$, $m$, $M_{off}$)
16:        end if
17:        $m\leftarrow m-M_{off}$
18:        $B\leftarrow B\cup$ SLO_AWARE_SCHEDULE($R_{off}$, $Q_{off}$, $t$, $c$, $m$)
19:        $Q_{recv}$.send($B$, $t$, $m$)
20:     end while
21:  end function
22:  function SCHEDULER
23:     Output: batched requests $B$
24:     scheduled requests $B$, latency budget $t$, memory budget $m$leftarrow $Q_{recv}$.get(block=True)
25:     if REQUEST_ARRIVAL then
26:        UPDATE_BUDGET($B$, $t$, $m$)
27:        if $t<0$ or $m<0$ then
28:           PREEMPT_UNTIL_FIT($B$, $t$, $m$)
29:        end if
30:     end if
31:     $Q_{send}$.send($B$)
32:     return $B$
33:  end function
A.2 SLO-Aware Prefix Sharing Maximization Algorithm
This section details the SLO-aware prefix sharing algorithm design in HyGen. For offline requests, we construct a prefix tree $T_{p}$ for prefix sharing, and a self-balancing BST $T_{f}$ for request freshness. A utility value between 0 and 1 is used to balance the chance between these two data structures. During scheduling, prefix sharing maximization algorithm retrieves waiting offline requests from either $T_{p}$ or $T_{f}$, and keeps their original order when scheduling running requests, effectively utilizing prefix sharing while avoiding starvation.
Algorithm 3 Prefix-sharing-aware offline scheduler
1:  Construct prefix tree $T_{p}$ and self-balanced BST $T_{f}$
2:  function PREFIX_SHARING_OFFLINE_SCHEDULE
3:     Input: running requests $R$, latency budget $t$, remaining chunk size $c$, memory budget $m$, utility value $u$
4:     Output: batched requests $B$
5:     for $r\in R$ do
6:        if $r$.state == DECODE then
7:           $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
8:           if $t>t_{req}$ then
9:              break
10:           end if
11:           $t\leftarrow t-t_{req}$, $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
12:        else if $r$.state == PREFILL then
13:           $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
14:           if $l>0$ then
15:              $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
16:           else
17:              break
18:           end if
19:        end if
20:     end for
21:     while $T_{p}$ is not empty do
22:        $rand\leftarrow$ RANDOM_NUMBER(0, 1)
23:        if $rand<u$ then
24:           $r\leftarrow T_{p}$.get_next_request()
25:        else
26:           $r\leftarrow T_{f}$.get_next_request()
27:        end if
28:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
29:        if $l>0$ then
30:           $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
31:           if $rand<u$ then
32:              $T_{f}$.remove_request($r$)
33:           else
34:              $T_{p}$.remove_request($r$)
35:           end if
36:        else
37:           break
38:        end if
39:     end while
40:     return $B$
41:  end function