### v2

# HyGen: Efficient LLM Serving via Elastic Online-Offline Request Co-location

Abstract
Large language models (LLMs) have facilitated a wide range of applications with distinct service-level objectives (SLOs), from latency-sensitive online tasks like interactive chatbots to throughput-oriented offline workloads like document summarization. The existing deployment model, which dedicates machines to each workload, simplifies SLO management but often leads to poor resource utilization.
This paper introduces HyGen, an interference-aware LLM serving system that enables efficient co-location of online and offline workloads while preserving latency requirements. HyGen incorporates two key innovations: (1) performance control mechanisms, including a latency predictor to estimate batch execution time and an SLO-aware profiler to quantify latency interference, and (2) SLO-aware offline scheduling policies that maximize serving throughput and prevent starvation, without compromising online serving latency. Our evaluation on production workloads shows that HyGen achieves up to 3.87$\times$ overall throughput and 5.84$\times$ offline throughput gains over online and hybrid serving baselines, respectively, while strictly satisfying latency SLOs.
1 Introduction
Large language models (LLMs) have emerged as transformative tools across diverse domains, handling both latency-critical online requests (e.g., chatbot interactions OpenAI (2022); Liu et al. (2024a)) and throughput-oriented offline tasks (e.g., document summarization Grand View Research (2023)). Online serving demands low and stable response times, measured by Time to First Token (TTFT) and Time Between Tokens (TBT), while offline tasks prioritize high throughput and resource utilization, often processing large batches with relaxed latency constraints. The disparity in these requirements has led most production deployments to segregate online and offline serving onto separate GPU clusters Sheng et al. (2023); Zhao et al. (2024); Patel et al. (2024); Stojkovic et al. (2024).
However, real-world LLM workloads demonstrate significant temporal variations in request load. Our analysis of production traces shows that in addition to the popular diurnal request arrival patterns, online request rates can vary by up to 3$\times$ within minutes (Section 3). To meet latency requirements under such bursty loads, service providers must provision GPU resources for peak demand, leading to substantial underutilization during off-peak hours.
This resource overprovisioning suggests an opportunity to improve resource efficiency via hybrid serving—co-locating online and offline workloads on the same GPU cluster. By opportunistically padding online requests with offline requests during periods of low online load, the system could maintain high GPU utilization while preserving latency guarantees for online requests. However, realizing this opportunity requires addressing several fundamental challenges.
First, LLM services exhibit diverse latency requirements across applications Liu et al. (2024a). Interactive chatbots require consistent response times with both low initial latency (TTFT) and smooth token generation (TBT), while batch processing tasks prioritize throughput over latency. These requirements often manifest in different statistical metrics—from strict P99 latency bounds to mean performance targets—making it difficult to establish unified resource-sharing policies.
Second, LLM workloads are inherently unpredictable in both their arrival patterns and resource demands. Request arrival rates exhibit both diurnal patterns and unpredictable short-term fluctuations. This variability is further complicated by uncertainty in resource demands—input sequences vary widely in length, and the number of output tokens can hardly be predicted until generation completes.
Third, co-locating online and offline workloads introduces interference. For instance, offline requests may delay time-sensitive online requests; using large batch sizes to improve offline request throughput can increase the latency of online requests. Effectively managing this interference while ensuring performance guarantees demands meticulous orchestration of resource sharing.
This paper presents HyGen, an interference-aware LLM serving system that elastically co-locates online and offline workloads. HyGen introduces several key techniques: (1) a latency predictor that accurately estimates the execution time of different request batches, (2) an interference-aware profiler that quantifies the performance interference of co-location, and (3) an adaptive scheduler that maximizes offline throughput while maintaining strict latency guarantees for online requests.
Our evaluation on production workloads shows that HyGen improves serving throughput by 3.87-5.84$\times$ over existing advances Agrawal et al. (2024), while guaranteeing strict SLO compliance. To summarize, this paper makes the following contributions:
  • We investigate the online request patterns in production LLM serving trace, revealing key challenges and opportunities in co-locating online and offline workloads.
  • We propose HyGen, a new serving system that enables efficient resource sharing, improving throughput with performance isolation.
  • We evaluate HyGen using production LLM serving traces to demonstrate HyGen’s effectiveness.
2 Related Work
LLM Online Serving.
Recent works have improved GPU computation through optimized kernels Dao et al. (2022); Ye et al. (2025), memory efficiency through efficient KV-cache management techniques Lin et al. (2024a); Liu et al. (2024b); Qin et al. (2024), distributed inference through various model parallelism strategies Shoeybi et al. (2019); Pope et al. (2023); Liu et al. (2023), and model compression approaches Xiao et al. (2023); Lin et al. (2024b); Dettmers et al. (2022). Orca Yu et al. (2022) introduces iteration-level scheduling to eliminate padding overhead in variable-length sequence batching and enables early exits for completed requests. vLLM Kwon et al. (2023) develops PagedAttention for efficient memory management, while Sarathi-Serve Agrawal et al. (2024) proposes chunked prefill to minimize inter-request interference within batches. However, few existing systems address the challenge of co-locating latency-critical online requests with offline workloads.
Throughput-Oriented LLM Inference.
As offline inference gains increasing traction, Huggingface Accelerate Hugging Face (2025), DeepSpeed ZeRO-Inference Aminabadi et al. (2022) and FlexGen Sheng et al. (2023) enable efficient LLM serving on commodity GPUs by offloading strategies. FlexGen aggregates memory from GPU, CPU and disk to maximize serving throughput. S3 Jin et al. (2023) utilizes output sequence length prediction technique, allowing more efficient memory allocation by avoiding over-allocation of the maximum sequence length.
Co-locating Workloads.
Co-locating latency-sensitive applications with batch applications has been widely adopted to improve resource utilization in datacenters Chen et al. (2019); Zhu & Erez (2016); Zhang et al. (2019); Iorgulescu et al. (2018); Patel & Tiwari (2020); Chen et al. (2023). Prior work has also explored resource multiplexing to serve multiple DNNs Dhakal et al. (2020); Tan et al. (2021); Zhang et al. (2023b), LLMs Duan et al. (2024), or LoRAs Chen et al. (2024); Sheng et al. (2024a); Wu et al. (2024). However, prior efforts have largely overlooked the opportunity to co-locate online and offline inference requests within the same LLM.
3 Background and Motivation
This section first introduces the core concepts of LLM serving systems (Section 3.1), then illustrates how these characteristics motivate our system designs (Section 3.2).
3.1 LLM Serving
LLM generation consists of two sequential stages: (1) Prefill Stage: It processes the input prompt in parallel to generate the first output token, computing and caching intermediate model representations (i.e., the KV cache) to avoid regeneration in subsequent tokens, and the (2) Decode Stage: the model generates tokens one at a time in an auto-regressive manner. Each iteration processes only the most recent token while utilizing the KV cache of previous tokens, reducing computational overhead.
Model executions in the prefill stage are compute-intensive due to the high degree of parallelism in processing the entire input sequence simultaneously. Yet in the decode phase, the same operations become memory-intensive due to the reliance on the KV cache Agrawal et al. (2024), and the total memory access cost grows linearly with the sequence length as each new token’s KV cache is appended.
Serving Workloads.
LLM serving deployment can be categorized into two primary scenarios. The first scenario is online serving, which targets real-time user interactions, such as chatbots, code assistants, and interactive applications OpenAI (2022); Microsoft (2023). This interactive nature often requires short Time to First Token (TTFT)—the latency from receiving a query to generating the first token—as well as the short Time Between Tokens (TBT), i.e., the inter-token generation latency in the decode stage.
The second scenario is offline serving, which prioritizes throughput over latency, measured in queries per second (QPS) or tokens per second (TPS). For example, OpenAI’s Batch API processes requests with relaxed latency requirements (up to 24 hours) at significantly lower costs compared to standard APIs OpenAI (2024). Applications include model benchmarking Hendrycks et al. (2021b); Meta-Team (2024), document processing Chen et al. (2021); Zhang et al. (2023a), data cleaning Narayan et al. (2022); Zheng et al. (2023), and synthetic data generation Allal et al. (2024).
3.2 Motivation and Challenges
(a) One-hour Request Rates
(b) Two-minute Request Rates
Figure 1: Request rate varies significantly in Microsoft Azure’s LLM service over one-hour and two-minute periods.
LLM serving systems face a critical resource utilization challenge due to the highly variable nature of their workloads. Figure 1 reports our analysis of Microsoft Azure’s LLM service Patel et al. (2024), showing that request rates can fluctuate dramatically—varying up to 3× within minutes while following broader diurnal patterns. This variability creates an inherent tension in resource provisioning: GPU clusters must be sized to handle peak loads, inevitably leading to resource underutilization during off-peak periods.
This load variability suggests an opportunity to improve GPU utilization by co-locating online requests with offline requests. During periods of low online traffic, the system could opportunistically schedule offline tasks to harvest idle GPU resources. While recent methodologies in simultaneous batching of prefill and decode requests set the premise for dynamic request co-location Agrawal et al. (2024), doing so at scale introduces several fundamental challenges:
  1. 1.
    Diverse Latency Requirements: Applications and even requests of an application have distinct latency requirements. For example, paid users require strict latency SLOs while free users accept more relaxed guarantees. How to flexiblely specify latency guarantees for online requests when sharing resources?
  2. 2.
    Massive Uncertainties: LLM serving faces temporal uncertainty—online requests arrive in unpredictable bursts with varying urgency levels—as well as resource demand uncertainty due to unpredictable output lengths. How to perform efficient scheduling in the wild?
  3. 3.
    Request Interference: Co-locating online and offline workloads introduces performance interference. Large batches of offline request can cause severe head-of-line blocking, delaying the processing of time-sensitive online requests. Worse, batching requests of long inputs with short interactive queries can elongate the latency of all requests in the batch by an order of magnitude Agrawal et al. (2024). How to account for interference in co-locating requests?
4 The HyGen Design
4.1 Overview
Figure 2: HyGen overview. Online requests and offline requests are asynchronously processed. Offline requests are opportunistically scheduled while retaining the batch latency budget.
In this paper, we present HyGen, a novel LLM serving system that enables efficient co-location of online and offline requests while maintaining latency guarantees. Unlike conventional LLM serving systems that segregate online and offline workloads onto separate GPU clusters, HyGen coordinates both types of requests through a two-phase scheduling approach, maximizing resource utilization during off-peak periods while preserving performance isolation.
Algorithm 1 HyGen SLO-aware scheduler
1:  function SLO_AWARE_SCHEDULE
2:     Input: running requests $R$, request queue $Q$,
3:      latency budget $t$, chunk size $c$, memory budget $m$
4:     Output: batched requests $B$
5:     $B\leftarrow\{\}$
6:     for $r\in R$.decode do
7:        $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
8:        if $t_{req}\leq t$ or PHASE ==absent$=$ = ONLINE then
9:           $t\leftarrow t-t_{req}$
10:           $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
11:        end if
12:     end for
13:     for $r\in R$.prefill $\cup Q$ do
14:        TRY_SCHEDULE:
15:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
16:        if $l>0$ then
17:           $t\leftarrow t-t_{req}$
18:           $c\leftarrow c-l$
19:           $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
20:           $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
21:        else
22:           if PHASE ==absent$=$ = ONLINE and $R\neq B$ then
23:              PERFORM_PREEMPTION($R$, $m$)
24:              goto TRY_SCHEDULE
25:           else
26:              break
27:           end if
28:        end if
29:     end for
30:     return $B$
31:  end function
Figure 2 illustrates the overall design of HyGen. To handle diverse SLOs and unpredictable workloads, HyGen maintains separate queues for online and offline requests, each managed by a dedicated request selector. This design makes HyGen compatible with various existing scheduling policies for different request types. Following Sarathi-Serve Agrawal et al. (2024), we employ stall-free batching and chunked-prefills, batching prefill and decode requests together for efficient utilization of compute resources and dynamic co-location of requests.
HyGen’s scheduling process occurs in two phases. The first (online) phase handles the online request queue to construct an initial batch, supporting various scheduling policies such as First-Come-First-Serve (FCFS) Kwon et al. (2023) and Virtual Token Counter (VTC) Sheng et al. (2024b). If the memory limit is reached and more online requests remain, a priority-based preemption mechanism is employed by prioritizing online requests over offline requests to minimize interference. Preempted offline requests are reclaimed into the offline queue, where they maintain their execution state to enable efficient rescheduling in subsequent batches.
In the second (offline) phase, the remaining capacity is filled with offline requests using a latency predictor to capture the latency impact of request filling. The predictor calculates the latency of the online batch and selects offline requests based on the available batch execution latency budget, i.e., the latency slack without violating the latency requirement of online requests. It predicts the latency cost for decode requests, and calculates a maximum chunked prefill length for prefill requests for the given latency budget. Offline requests are scheduled in order until reaching the resource capacity, based on our SLO-aware scheduling policy. A formalized algorithm of the SLO-aware scheduling process is displayed in Algorithm 1. Appendix A.1 further presents a holistic workflow of two-phase scheduling, including the management of online and offline request queues and implementation of our SLO-aware scheduling algorithm.
Section 4.2 describes our performance control mechanisms, including the latency predictor and SLO-aware profiler that enable accurate resource allocation. Section 4.3 presents our SLO-aware offline scheduling policies that maximize throughput while maintaining SLOs.
4.2 Performance Control Mechanisms
Co-locating online and offline requests requires precise control over execution timing and resource allocation. The key challenge lies in accurately estimating the latency impact of scheduling decisions to ensure end-to-end SLO compliance. This section introduces a latency predictor for estimating batch execution latency and an SLO-aware profiler for translating the estimates into scheduling decisions.
Latency Predictor.
The design of our latency predictor is guided by three key requirements. First, it must provide fast inference to support real-time scheduling decisions. Second, it needs to be robust across varying workload patterns to maintain reliable performance. Third, it should be adaptable to different hardware configurations to accurately capture their unique performance characteristics.
The execution time of an LLM serving batch is primarily determined by two distinct processing stages with different computational patterns. The prefill stage exhibits quadratic complexity due to attention computations, with latency growing quadratically with input sequence length. The total load of this stage depends on both the number of requests and their individual sequence lengths. In contrast, the decode stage shows linear scaling, with computational requirements growing proportionally with the number of tokens. We can model the batch execution time below to capture these characteristics:
$T_{batch}=f(S_{p},S_{d},S_{p}^{2},S_{d}^{2},N_{p},N_{d})$ (1)
where $S_{p}$ and $S_{d}$ represent the total number of prefill and decode tokens in the batch, respectively. The quadratic terms ($S_{p}^{2}$ and $S_{d}^{2}$) account for non-linear scaling effects, particularly in the prefill phase. $N_{p}$ and $N_{d}$ represent the number of prefill and decode requests in the batch, respectively.
We employ linear regression as the prediction model because of its efficiency and effectiveness. Training data for the model is collected by systematically profiling target hardware across diverse batch compositions, varying in the number of requests in different phases, sequence length distributions, and total batch sizes. The linear model enables rapid evaluation of different batch compositions during scheduling, while its simple feature set ensures stable predictions across varying conditions.
SLO-Aware Profiling.
The latency predictor provides accurate latency estimates for filling offline requests. Our SLO-aware profiler leverages a latency budget to ensure SLO compliance in scheduling.
The profiler first analyzes the given combination of workload and SLO to establish viable latency budget ranges. Given that a larger batch size and longer inputs will increase batch latency, the profiler test-runs latency budgets within the range to check their compliance with the given SLO and employs binary search to decide an upper limit that meets the overall SLO for online requests. During deployment, this latency budget is then used as the batch latency limit in the two-phase SLO-aware scheduling process (Section 4.1) to ensure SLO compliance. This profiling enables three key capabilities. First, it determines appropriate latency thresholds that maintain SLO compliance for different workload types. Second, it provides flexible adaptation by adjusting budgets based on changing workload characteristics and performance requirements. Third, it establishes a robust foundation for hybrid scheduling by accounting for both online and offline workload patterns.
4.3 SLO-Aware Offline Scheduling Policies
After scheduling online requests, our offline scheduling policy repurposes the residual capacity to maximize throughput while ensuring fairness Sheng et al. (2024b).
To further optimize the serving throughput for offline requests, HyGen employs an SLO-aware Prefix Sharing Maximization (PSM) strategy. Prefix sharing is a widely-adopted technique for reusing the KV cache of shared input prefix between requests Gim et al. (2024); Zheng et al. (2024a); Zhao et al. (2024); Zheng et al. (2024b).
Figure 3: HyGen respects latency requirements in co-locating requests.
Prefix Sharing Maximization Strategy.
Our Prefix Sharing Maximization (PSM) strategy organizes offline requests into a prefix tree that follows the structure of a Trie tree, capturing prefix sharing characteristics of all offline requests, with each leaf node representing an offline request. The priority of each request is determined by the Depth-First Search (DFS) order of the prefix tree, where requests with the greatest prefix sharing potential are scheduled together, maximizing prefix cache reuse among requests. Subsequently, HyGen performs SLO-aware offline scheduling (Section 4.1) using this order to maximize prefix sharing, reducing redundant computation and improving throughput. Additionally, the prefix tree structure ensures fast insertion of new arrivals and deletion of scheduled requests, guaranteeing efficient scheduling during runtime. A formalized algorithm is demonstrated in Appendix A.2. For example, consider a system that can process two offline requests per batch with the following request queue: (What is ML, How to code, What is AI, How to debug). Under traditional FCFS scheduling, requests are processed in arrival order: (What is ML, How to code), (What is AI, How to debug), resulting in no prefix sharing opportunities. In contrast, PSM’s prefix-aware scheduling reorders requests as: (What is ML, What is AI), (How to code, How to debug), enabling KV-cache reuse through shared prefixes.
The PSM strategy demonstrates strong extendability. Under certain scenarios, the vanilla PSM strategy may lead to starvation for requests with minimal prefix-sharing potential, blocked by continuous arrivals for an extended period. To illustrate this challenge, consider a request queue: (What is ML, What is AI, How to code, What is DL). When new requests arrive with similar prefixes (What is LLM, What is DNN), a naive prefix-sharing policy would continuously prioritize requests sharing the What is prefix, potentially starving the How to code request indefinitely. This issue can be mitigated by an extended version of our PSM policy, combining maximum prefix sharing with request freshness by using a utility ratio to ensure a balance between efficiency and fairness. Based on the utility ratio, a new offline request would be selected from the DFS order of the prefix tree or the most stale request from a self-balanced binary search tree sorted by freshness. A detailed algorithm is in Appendix A.3. These enhancements can make the PSM strategy more practical for real-world deployments, retaining its efficiency while improving fairness and adaptability.
5 Performance Evaluation
Figure 4: HyGen improves serving throughput under varying SLOs.
5.1 Evaluation Setup
Implementation and Testbeds.
We implement HyGen on top of vLLM Kwon et al. (2023); vLLM Project (2023) and Sarathi-Serve Agrawal et al. (2024); Sarathi-Serve Project (2024), with 1,300 lines of additional code. We evaluate HyGen on two server configurations: one with 4 NVIDIA A100 GPUs (40GB VRAM each) and the other with 4 NVIDIA A40 GPUs (48GB VRAM each). Both servers have 64 CPU cores, 256GB DDR4 RAM, and a 1.5TB NVMe SSD.
Models and Workloads.
We evaluated HyGen using Llama2-7B Touvron et al. (2023) and Qwen-14B Bai et al. (2023) models on A100 and A40 GPUs, respectively.
For the online workload, we used the conversation trace from Azure LLM inference trace 2023 Patel et al. (2024), a one-hour production trace with real-world requests and timestamps. We randomly sampled the trace to achieve the desired QPS that suits our hardware serving capacity. Specifically, within a time duration of $T$ seconds, we would sample $T\times Q$ requests to suit a desired QPS $Q$. For the offline workload, we used arXiv summarization Cohan et al. (2018), a dataset for long document summarization.
Metrics.
For interference evaluation, we focus on both Time to First Token (TTFT) and Time Between Tokens (TBT) and their respective mean and 99th percentile (P99) values. We measure throughput using tokens per second (TPS) and queries per second (QPS).
Baselines.
For pure online inference, we use Sarathi Agrawal et al. (2024) as our baseline. For hybrid serving, we compare HyGen with two baselines for interference and throughput evaluation, respectively: (1) Sarathi++: We implement our online-first scheduling policy on Sarathi to support hybrid serving, including the request management and preemption handling policies introduced in Section 4.1. (2) HyGen*: To evaluate the throughput benefit of our HyGen design, we further improve Sarathi++ to an SLO-aware serving system, HyGen*. Besides inheriting the serving policies from Sarathi++, HyGen* serves offline requests at a specific offline QPS to control overall interference. The offline QPS is profiled using a similar design with the HyGen profiler to guarantee bounded SLO interference.
Figure 5: HyGen latency predictor achieves high accuracy for batch latency prediction.
Figure 6: Prefix Sharing Maximization policy achieves higher serving capacity.
Figure 7: SLO-aware profiler bridges the gap between batch latency and overall SLO.
Figure 8: Hygen dynamically controls offline throughput according to online workload and latency budget.
(a) SLOs
(b) Throughput
Figure 9: Hygen meets SLOs and achieves higher throughput for Yi-34B model using TP=2, PP=2.
Figure 10: HyGen meets SLO under various online QPS settings.
Figure 11: HyGen is able to meet multiple SLOs simultaneously.
(a) SLOs
(b) Throughput
Figure 12: Hygen meets SLOs and achieves higher throughput for Mooncake trace and Mistral-7B model.
(a) SLOs
(b) Throughput
Figure 13: Hygen meets SLOs and achieves higher throughput.
Figure 14: HyGen is robust to predictor accuracy.
Figure 15: Hygen dynamically adjusts offline throughput.
5.2 End-to-end Performance
HyGen respects latency requirements in co-locating requests. We tested HyGen on four SLO metrics, mean TBT, P99 TBT, mean TTFT, and P99 TTFT. For each metric, we set varying interference tolerance ratios as the SLO. Figure 3 shows that HyGen is able to control the interference and guarantees to meet specific SLOs across our settings. Compared with Sarathi++, an SLO-unaware system that yields the same result for all metrics and tolerance ratios, HyGen showcases efficient SLO-aware latency control.
HyGen improves serving throughput. Figure 4 shows the offline throughput of HyGen for various metrics and tolerance ratios. Through efficient request co-location, HyGen improves overall serving throughput by up to 3.87$\times$ compared to pure online serving. Under the same SLO, HyGen consistently achieves higher throughput compared to HyGen*, yielding up to 5.84×5.84$\times$5.84 × offline throughput gain. This verifies that our fine-grained latency predictor and SLO-aware profiler designs achieve higher serving efficiency compared to their simplified counterparts.
5.3 Performance Breakdown
Accuracy of latency predictor.
We evaluate the accuracy of our latency predictor on Llama2-7B and Qwen-14B using Azure LLM Inference trace mixed with arXiv summarization dataset. Figure 5 shows that our latency predictor achieves a mean absolute percentage error of only 1.78% and 1.07%, confirming its high accuracy.
Impact of prefix sharing.
To test HyGen’s compatibility with prefix sharing, we conducted a simulation experiment using Azure LLM Inference as the online trace and MMLU Hendrycks et al. (2021a; b) as the offline dataset on a Llama2-7B model. In our simulation, we deducted the shared prompt prefix length between consecutive offline requests to simulate prefix sharing. Figure 7 shows that HyGen yields up to 4×4$\times$4 × offline throughput gain with its prefix sharing maximization scheduling policy.
Impact of SLO-aware profiler.
To demonstrate the effect of HyGen’s SLO-aware profiler, we compared it with a simple strategy that sets the desired mean TBT SLO as the batch latency budget. Figure 7 shows the performance gap between individual batch latencies and overall mean TBT, illustrating how the SLO-aware profiler bridges this gap for controlled SLO in hybrid serving.
Breakdown by time.
Figure 8 shows a temporal throughput breakdown of HyGen. At runtime, HyGen dynamically adjusts offline throughput based on online workload and overall latency budget, batching offline requests more aggressively during online QPS troughs and reducing offline throughput during online bursts, harnessing compute resources in an adaptive manner.
5.4 Ablation Studies
Impact of model parallelisms.
To evaluate HyGen’s effectiveness in distributed inference, we deployed the Yi-34B model Young et al. (2024) on a server with 4 NVIDIA A40 GPUs using tensor-parallelism (TP) Shoeybi et al. (2019) and pipeline-parallelism (PP) Athlur et al. (2022); Narayanan et al. (2019); Huang et al. (2019) with degree 2 for each dimension. Using Azure LLM Inference and arXiv summarization workloads, Figure 9 shows that HyGen maintains its ability to meet SLOs and achieves higher offline throughput (up to 1.89×1.89$\times$1.89 ×) than the baseline.
Impact of SLO requirements.
We further evaluate HyGen’s ability to meet stringent SLOs under varying online QPS settings on the four aforementioned metrics, each with 5% interference tolerance. Figure 10 shows that HyGen meets stringent SLOs for all metrics. We further demonstrate HyGen’s ability to meet multiple SLOs at the same time. By testing HyGen with a fixed P99 TTFT interference ratio (8%) and mean TBT interference ratios ranging from 10% to 50%, Figure 11 shows that at a lower mean TBT tolerance, HyGen’s performance is bounded by mean TBT SLOs; after reaching the fixed P99 TTFT SLO, mean TBT stops increasing in order to keep P99 TTFT under control.
Impact of models and datasets.
We further evaluate HyGen’s adaptability on two more experiments: The first using Mistral-7B model with Mooncake trace Qin et al. (2024), a trace containing request lengths and timestamps taken from real-world servers, as the online trace, and arXiv summarization as the offline trace; The second experiment uses Llama2-7B model with Azure LLM Inference trace and replaced the offline dataset with CNN/DailyMail summarization dataset See et al. (2017); Hermann et al. (2015). An analysis of the Mooncake trace is shown in Appendix B. Figure 12 and Figure 13 show that HyGen achieves superior performance than its counterparts under these settings.
Impact of predictor accuracy.
We tested HyGen’s robustness using several pre-trained LR latency predictors with varying prediction accuracy taken from other workloads and tested them on Azure LLM Inference trace and arXiv summarization dataset. Figure 15 shows how predictor accuracy (measured in mean absolute percentage error) affects offline throughput under the same P99 TBT SLO. HyGen remains robust across different accuracy settings. Additionally, our LR-based latency predictor is lightweight for training, with only $\sim$15ms training time for over 80,000 samples on CPUs.
Impact of online arrival rate.
Under a high online arrival rate, online requests may approach system serving capacity, limiting the headroom for request co-location. Figure 15 shows the effect of online QPS on offline throughput with 5% P99 TBT tolerance. As online load increases, HyGen adjusts offline throughput based on the system’s residual serving capacity while maintaining higher throughputs.
6 Conclusion
This paper introduces HyGen, an LLM serving system that enables efficient co-location of online and offline workloads. We employ control mechanisms to predict and manage interference impacts, and a scheduling policy to opportunistically schedule offline serving. Evaluation on production workloads demonstrates that HyGen improves serving throughput by 3.87-5.84$\times$ while maintaining strict latency SLOs.
Appendix
Appendix A Algorithms
A.1 HyGen Two-phase Scheduling Algorithm
This section gives a detailed and formalized demonstration of the two-phase scheduling algorithm introduced in Section 4.1. In each scheduling step, the workflow invokes the SLO-aware scheduling process in Algorithm 1 twice (line 13 and line 18) to form a hybrid batch with online and offline request co-location while respecting latency and memory limits. To reduce the scheduling overhead of HyGen, we employ a message queue for asynchronous communication between the main process and the offline scheduler. After each scheduling step, the main process sends the metadata of batched requests to the message queue. The offline scheduler first calculates the expected status of each request based on scheduling decisions from the previous batch, and then runs a scheduling simulation to generate offline request scheduling decisions using our latency predictor and the profiled latency budget. The offline scheduling decisions are then sent back to the main process using the message queue and used for the next scheduling step. To support pipeline parallelization, a scheduling history archive of $K$ steps is kept by the offline scheduler for pipeline parallelization degree $K$, in order to have a holistic view of every request running in each pipeline stage at the time.
Algorithm 2 HyGen two-phase scheduler
1:  global $Q_{on}$ (online request queue), $Q_{off}$ (offline request queue)
2:  global $R_{on}$ (online request running list), $R_{off}$ (offline request running list)
3:  global $Q_{send}$, $Q_{recv}$ (message queues)
4:  global $L$ (latency budget), $M$ (memory budget), $C$ (chunk size), $M_{off}$ (offline memory)
5:  function ASYNC_SCHEDULER
6:     while True do
7:        scheduled requests $S\leftarrow Q_{send}$.get(block=True)
8:        UPDATE_REQUEST_STATUS($S$, $R_{on}$, $R_{off}$)
9:        batched requests $B\leftarrow$}
10:        latency budget $t\leftarrow L$
11:        memory budget $m\leftarrow$ GET_FREE_MEMORY() + $M_{off}$
12:        chunk size $c\leftarrow C$
13:        $B\leftarrow B\cup$ SLO_AWARE_SCHEDULE($R_{on}$, $Q_{on}$, $t$, $c$, $m$)
14:        if $m<M_{off}$ then
15:           PREEMPT_OFFLINE($R_{off}$, $m$, $M_{off}$)
16:        end if
17:        $m\leftarrow m-M_{off}$
18:        $B\leftarrow B\cup$ SLO_AWARE_SCHEDULE($R_{off}$, $Q_{off}$, $t$, $c$, $m$)
19:        $Q_{recv}$.send($B$, $t$, $m$)
20:     end while
21:  end function
22:  function SCHEDULER
23:     Output: batched requests $B$
24:     scheduled requests $B$, latency budget $t$, memory budget $m$leftarrow $Q_{recv}$.get(block=True)
25:     if REQUEST_ARRIVAL then
26:        UPDATE_BUDGET($B$, $t$, $m$)
27:        if $t<0$ or $m<0$ then
28:           PREEMPT_UNTIL_FIT($B$, $t$, $m$)
29:        end if
30:     end if
31:     $Q_{send}$.send($B$)
32:     return $B$
33:  end function
A.2 SLO-Aware Prefix Sharing Maximization Algorithm
This section details the SLO-aware prefix sharing maximization algorithm design in HyGen. We construct a prefix tree $T_{p}$ to capture prefix sharing charateristics among offline requests. During scheduling, offline requests are selected in the DFS order of the prefix tree, and deleted once being scheduled. Additionally, running requests keep their original DFS order in future scheduling process, effectively utilizing prefix sharing.
Algorithm 3 Prefix-sharing-aware offline scheduler
1:  Construct prefix tree $T_{p}$
2:  function PREFIX_SHARING_OFFLINE_SCHEDULE
3:     Input: running requests $R$, latency budget $t$, remaining chunk size $c$, memory budget $m$
4:     Output: batched requests $B$
5:     $B\leftarrow\{\}$
6:     for $r\in R$ do
7:        if $r$.state == DECODE then
8:           $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
9:           if $t>t_{req}$ then
10:              break
11:           end if
12:           $t\leftarrow t-t_{req}$
13:           $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
14:        else if $r$.state == PREFILL then
15:           $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
16:           if $l>0$ then
17:              $t\leftarrow t-t_{req}$
18:              $c\leftarrow c-l$
19:              $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
20:              $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
21:           else
22:              break
23:           end if
24:        end if
25:     end for
26:     while $T_{p}$ is not empty do
27:        $r\leftarrow T_{p}$.get_next_request()
28:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
29:        if $l>0$ then
30:           $t\leftarrow t-t_{req}$
31:           $c\leftarrow c-l$
32:           $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
33:           $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
34:           $T_{p}$.remove_request($r$)
35:        else
36:           break
37:        end if
38:     end while
39:     return $B$
40:  end function
A.3 Extended SLO-Aware Prefix Sharing Maximization Algorithm
This section details the extended version of the SLO-aware prefix sharing maximization algorithm, enhanced with fairness-aware scheduling. For offline requests, we construct a prefix tree $T_{p}$ for prefix sharing, and a self-balancing BST $T_{f}$ for request freshness. A utility ratio between 0 and 1 is used to balance the chance between these two data structures. During scheduling, the extended prefix sharing maximization algorithm retrieves waiting offline requests from either $T_{p}$ or $T_{f}$, based on the utility ratio. The selected request is then deleted from both data structures to ensure synchronization. This solution balances prefix sharing and fairness without disrupting the DFS order of the prefix tree while avoiding possible starvation.
Algorithm 4 Prefix-sharing-aware offline scheduler
1:  Construct prefix tree $T_{p}$ and self-balanced BST $T_{f}$
2:  function PREFIX_SHARING_OFFLINE_SCHEDULE
3:     Input: running requests $R$, latency budget $t$, remaining chunk size $c$, memory budget $m$, utility value $u$
4:     Output: batched requests $B$
5:     $B\leftarrow\{\}$
6:     for $r\in R$ do
7:        if $r$.state == DECODE then
8:           $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
9:           if $t>t_{req}$ then
10:              break
11:           end if
12:           $t\leftarrow t-t_{req}$, $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
13:        else if $r$.state == PREFILL then
14:           $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
15:           if $l>0$ then
16:              $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
17:           else
18:              break
19:           end if
20:        end if
21:     end for
22:     while $T_{p}$ is not empty do
23:        $rand\leftarrow$ RANDOM_NUMBER(0, 1)
24:        if $rand<u$ then
25:           $r\leftarrow T_{p}$.get_next_request()
26:        else
27:           $r\leftarrow T_{f}$.get_next_request()
28:        end if
29:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
30:        if $l>0$ then
31:           $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
32:           $T_{p}$.remove_request($r$)
33:           $T_{f}$.remove_request($r$)
34:        else
35:           break
36:        end if
37:     end while
38:     return $B$
39:  end function
Appendix B Mooncake Trace Analysis
This section provides an analysis of Mooncake trace, a production trace taken from real-world server with request lengths and request arrival timestamps. Figure 16 shows the varying request arrival rates of Mooncake trace over one-hour and ten-minute periods. This analysis further demonstrates the fluctuant characteristic of LLM services.
(a) 1 hour
(b) 10 minutes
Figure 16: Request rate varies in Moonshot Mooncake’s LLM service over one-hour and ten-minute periods.