### v4

# HyGen: Efficient LLM Serving via Elastic Online-Offline Request Co-location

Abstract
Large language models (LLMs) have facilitated a wide range of applications with distinct service-level objectives (SLOs), from latency-sensitive online tasks like interactive chatbots to throughput-oriented offline workloads like data synthesis. The existing deployment model, which dedicates machines to each workload, simplifies SLO management but often leads to poor resource utilization.
[…]
1 Introduction
Large language models (LLMs) have emerged as transformative tools across diverse domains, handling both latency-critical online requests (e.g., chatbot interactions andes ; chatgpt ) and throughput-oriented offline tasks (e.g., document summarization llm-market-report ). Online serving demands low and stable response times, measured by Time to First Token (TTFT) and Time Between Tokens (TBT), while offline tasks prioritize high throughput and resource utilization, often processing large batches with relaxed latency constraints. The disparity in these requirements has led most production deployments to segregate online and offline serving onto separate clusters to avoid interference patel2024splitwise ; sheng2023flexgen ; stojkovic2024dynamollm ; zhao2024blendserve .
[…]
This resource overprovisioning suggests an opportunity to improve resource efficiency via hybrid serving—co-locating online and offline workloads on the same inference engine instance. By opportunistically padding online requests with offline requests during periods of low online load, the system could maintain high GPU utilization while preserving latency guarantees for online requests. However, realizing this opportunity requires addressing several fundamental challenges.
First, LLM services exhibit diverse latency requirements across applications andes . Interactive chatbots require consistent response times with both low initial latency (TTFT) and smooth token generation (TBT), while batch processing tasks prioritize throughput over latency. These requirements often manifest in different statistical metrics—from strict P99 latency bounds to mean performance targets—making it difficult to establish unified resource-sharing policies.
Second, LLM workloads are inherently unpredictable in both their arrival patterns and resource demands. Request arrival rates exhibit both diurnal patterns and unpredictable short-term fluctuations. This variability is further complicated by uncertainty in resource demands—input sequences vary widely in length, and the number of output tokens can hardly be predicted until generation completes.
Third, co-locating online and offline workloads introduces interference. For instance, offline requests may delay time-sensitive online requests; using large batch sizes to improve the throughput of offline requests can increase the latency of online requests. Effectively managing this interference while ensuring latency guarantees demands meticulous orchestration of resource sharing.
This paper presents HyGen, an interference-aware LLM serving system that elastically co-locates online and offline workloads. HyGen introduces several key techniques: (1) a latency predictor that accurately estimates the execution time of different request batches, (2) an interference-aware profiler that quantifies the performance interference of co-location, and (3) an adaptive scheduler that maximizes offline throughput while maintaining strict latency guarantees for online requests.
[…]
  • Key insights on the feasibility and benefits of co-locating online and offline LLM workloads, derived from systematic characterization of production traces.
  • A statistical latency prediction model that accurately captures the relationship between batch composition and execution latency, accounting for quadratic complexity in prefill and linear scaling in decode phases.
  • A novel scheduling system that dynamically co-locates online and offline workloads, formulated as a constrained optimization problem that maximizes throughput with prefix sharing while preserving strict latency SLOs and providing theoretical guarantees for fairness.
[…]
2 Related Work
LLM Inference Optimization.
Recent advances in LLM inference through kernel optimization flashattention ; flashinfer , compilation frameworks ma2025intelligen ; wu2025mirage ; zheng2023einnet , and scheduling algorithms lin2024infinite ; lin2024parrot ; sun-etal-2025-disco have substantially improved LLM serving performance. Among these, a particularly important development is predictive scheduling, which aims to estimate request difficulty ding2024hybrid ; ong2024routellm or generation length fu2024efficient ; hu2024inference ; jin2023s ; patke2024queue ; qiu2024power . Contrasting these efforts, HyGen uniquely predicts batch execution time, providing precise control over request interference during co-located workloads. Additionally, the concept of request prefix sharing has been explored by gim2024prompt ; juravsky2024hydragen ; zheng2024sglang , where shared prefixes across requests optimize resource use. HyGen takes this a step further by applying a prefix sharing maximization strategy to opportunistically schedule offline requests, utilizing residual capacity from a primary, SLO-bound online workload. Moreover, HyGen introduces a fairness-aware extension to the standard prefix-sharing maximization (PSM) method, addressing the issue of starvation often encountered in naive designs.
In parallel, the increasing demand for offline LLM inference has driven two key trends. First, platforms like Huggingface Accelerate hugging_face_accelerate , DeepSpeed ZeRO-Inference aminabadi2022deepspeed , and FlexGen sheng2023flexgen enable efficient inference on commodity hardware through memory offloading across GPUs, CPUs, and disk. Second, cloud providers, including OpenAI’s Batch API openai2024batchapi , have launched specialized services optimized for throughput rather than latency, processing millions of requests daily for large-scale data processing tasks.
Workloads Co-location.
In data center environments, co-locating latency-sensitive applications with batch applications has been explored to improve resource utilization chen2023olpart ; chen2019parties ; patel2020clite ; tempo-arixv25 . However, these approaches generally overlook the unique characteristics of LLM inference. In the context of LLM inference, several works have investigated the co-location of various model types and tasks. For instance, Punica chen2024punica , S-LORA sheng2024slora , and dLoRA wu2024dlora batch requests from different LoRAs, while MuxServe duanmuxserve multiplexes resources across multiple LLMs. In contrast, HyGen focuses on batching online and offline requests, addressing distinct optimization problems that are not interchangeable.
3 Background and Motivation
This section first introduces the background of LLM serving deployment (Section 3.1), then illustrates how these characteristics motivate our system designs (Section 3.2).
3.1 LLM Serving
Large-scale inference clusters consist of multiple serving instances, with a router intelligently directing incoming requests to the most suitable instances srivatsa2024preble ; iccache-sosp25 . Each instance, which could be based on architectures like vLLM kwon2023efficient or SGLang zheng2024sglang , typically employs iteration-level scheduling yu2022orca and chunked prefill sarathi . This setup enables decode requests to perform an additional decoding step, while prefill requests are limited to a fixed token budget. These two types of requests are processed together within a single iteration, optimizing resource utilization.
LLM serving deployment can be categorized into online serving and offline serving scenarios. Online serving targets real-time user interactions, such as chatbots, code assistants, and interactive applications githubGitHubCopilot ; chatgpt . This interactive nature often requires a short Time to First Token (TTFT) as well as a short Time Between Tokens (TBT). Offline serving prioritizes throughput over latency, measured in queries per second (QPS) or tokens per second (TPS). For example, OpenAI’s Batch API processes requests with relaxed latency requirements (up to 24 hours) at significantly lower costs compared to standard APIs openai2024batchapi . Applications include model benchmarking hendryckstest2021 ; dubey2024llama3herdmodels , document processing chen2021spreadsheetcoder ; zhang-etal-2023-summit , data cleaning narayan2022foundationmodelswrangledata ; zheng2023judgingllmasajudgemtbenchchatbot , and data synthesis huggingfaceCosmopediaCreate .
3.2 Motivation and Challenges
Figure 1: Request rate varies significantly in Microsoft Azure’s LLM service over one-hour and two-minute periods.
Figure 2: HyGen Overview. Online and offline requests are processed asynchronously, with offline requests opportunistically scheduled to respect latency budgets.
LLM serving systems face a critical resource utilization challenge due to the highly variable nature of their workloads. Figure 2 reports our analysis of Microsoft Azure’s LLM service patel2024splitwise , showing that request rates can fluctuate dramatically—varying up to 3× within minutes while following broader diurnal patterns. This variability creates an inherent tension in resource provisioning: serving clusters must be sized to handle peak loads, leading to resource underutilization during off-peak periods.
This load variability suggests an opportunity to improve resource utilization by co-locating online requests with offline requests. For each serving instance in the cluster, during periods of low online traffic, it can opportunistically schedule offline tasks to harvest idle resources. While recent methodologies in simultaneous batching of prefill and decode requests set the premise for dynamic request co-location sarathi , doing so at scale introduces several fundamental challenges:
  1. 1.
    Diverse Latency Requirements: Applications and even requests of an application have distinct latency requirements. For example, paid users require strict latency SLOs while free users accept more relaxed guarantees. How to respect diverse latency requirements in flight?
  2. 2.
    Massive Uncertainties: LLM serving faces temporal uncertainty—online requests arrive in unpredictable bursts with varying urgency levels—as well as resource demand uncertainty due to unpredictable output lengths. How to perform efficient scheduling in the wild?
  3. 3.
    Request Interference: Co-locating online and offline workloads introduces performance interference. Large batches of offline requests can cause severe head-of-line blocking, delaying the processing of time-sensitive online requests. Worse, batching requests of long inputs with short interactive queries can elongate the latency of all requests in the batch by an order of magnitude sarathi . How to account for interference in co-locating requests?
4 The HyGen Design
4.1 Overview
HyGen introduces a novel approach to integrate online and offline requests while maintaining strict latency guarantees. As shown in Figure 2, HyGen employs a dual-queue architecture that separates latency-sensitive and throughput-oriented requests. This design accommodates diverse SLOs and variable workloads while remaining compatible with existing scheduling policies within each queue. We note that HyGen functions as an instance-level scheduler, receiving requests from an upstream system-level router (e.g., Preble srivatsa2024preble ). As a result, both the request concurrency and scheduling overhead at each instance are inherently bounded. HyGen’s two-phase scheduling operates as follows (see Appendix A.1 for the asynchronous two-queue workflow and message passing details):
  • The online phase prioritizes latency-sensitive requests, forming an initial batch using established policies such as First-Come-First-Serve (FCFS) kwon2023efficient or fairness request scheduling sheng2024fairness . We introduce a priority-based preemption mechanism that protects online request performance by selectively preempting offline requests. Currently, HyGen preserves execution state for preempted requests, while its architecture supports various preemption mechanisms—including state discarding, preservation, and swapping—as categorized by InferCept abhyankar2024infercept .
  • The offline phase uses our latency predictor to allocate remaining capacity to throughput-oriented requests. This predictor accurately estimates the latency impact of each potential offline request addition, determining either decode request latency costs or maximum chunked prefill lengths that fit within the available latency budget without violating online SLOs.
Algorithm 1 formalizes our scheduling approach, with implementation details and complexity analysis ($O(n)$ where $n$ is the number of requests) provided in Appendices A.1 and A.4. Appendix C presents a novel cluster serving paradigm that addresses the longstanding tradeoff between SLO compliance and resource utilization based on HyGen.
Algorithm 1 HyGen SLO-aware scheduler
1:  function SLO_AWARE_SCHEDULE
2:  Input: running requests $R$, request queue $Q$,
3:     latency budget $t$, chunk size $c$, memory budget $m$
4:  Output: batched requests $B$
5:  $B\leftarrow\{\}$
6:  for $r\in R$.decode do
7:   $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)    // predict latency of the decoding request
8:   if $t_{req}\leq t$ or PHASE ==== ONLINE then
9:    // schedule request if it is: 1. online, or 2. offline and enough latency budget left
10:    $t\leftarrow t-t_{req}$
11:    $B\leftarrow B\cup$ {($r$, 0, $t_{req}$)}
12:  for $r\in R$.prefill $\cup Q$ do
13:   TRY_SCHEDULE:    // try to schedule a prefilling or waiting request
14:   // get the max number of tokens allowed under memory and latency budget
15:   $l,t_{req}\leftarrow$ PREDICTOR.get_max_tokens($t$, $c$, $m$, $r$)
16:   if $l>0$ then
17:    // schedule request
18:    $t\leftarrow t-t_{req}$
19:    $c\leftarrow c-l$
20:    $m\leftarrow m$- GET_NUM_BLOCKS($l$)
21:    $B\leftarrow B\cup$ {($r$, $l$, $t_{req}$)}
22:   else
23:    if PHASE ==== ONLINE and $R\neq B$ then
24:     PERFORM_PREEMPTION($R$, $m$)    // preempt request with lower priority
25:     goto TRY_SCHEDULE    // try to schedule again
26:    else
27:     break
28:  return $B$
4.2 Performance Control Mechanisms
Co-locating online and offline requests under diverse SLO requirements requires precise control over resource allocation. The key challenge lies in accurately estimating the latency impact of scheduling decisions to ensure SLO compliance. This section introduces a latency predictor for estimating batch execution latency and an SLO-aware profiler for translating the estimates into scheduling decisions.
Latency Predictor.
The design of our latency predictor is guided by three key requirements. First, it must provide fast inference to support real-time scheduling decisions. Second, it needs to be robust across varying workload patterns to maintain reliable performance. Third, it should be adaptable to different hardware configurations to accurately capture their unique performance characteristics.
The execution time of an LLM serving batch is primarily determined by two distinct processing stages with different computational patterns. The prefill stage exhibits quadratic complexity due to attention computations, with latency growing quadratically with input sequence length. The total load of this stage depends on both the number of requests and their individual sequence lengths. In contrast, the decode stage shows linear scaling, with computational requirements growing proportionally with the number of tokens. We can model the batch execution time below to capture these characteristics:
$T_{batch}=f(S_{p},S_{d},S_{p}^{2},S_{d}^{2},N_{p},N_{d})$ (1)
where $S_{p}$ and $S_{d}$ represent the total number of prefill and decode tokens in the batch, respectively. The quadratic terms ($S_{p}^{2}$ and $S_{d}^{2}$) account for non-linear scaling effects, particularly in the prefill phase. $N_{p}$ and $N_{d}$ represent the number of prefill and decode requests in the batch, respectively.
We employ linear regression as the prediction model because of its efficiency and effectiveness. Training data for the model is collected by systematically profiling target hardware across diverse batch compositions, varying in the number of requests in different phases, sequence length distributions, and total batch sizes. The linear model enables rapid evaluation of varying batch compositions during scheduling, while its simple feature set ensures stable predictions across varying conditions. Further discussion on the implementation and expandability of the LR predictor can be found in Appendix B.
SLO-aware Profiling.
The latency predictor provides accurate latency estimates for filling offline requests. Our SLO-aware profiler leverages a latency budget to ensure SLO compliance in scheduling. The profiler first analyzes the given combination of workload and SLO to establish viable latency budget ranges. Given that larger batch sizes and longer inputs will increase latency, the profiler test-runs latency budgets within the range to check their compliance with the given SLO and employs binary search to decide an upper limit that meets the overall SLO for online requests. During deployment, this latency budget is used as the batch latency limit in the two-phase scheduling process (Section 4.1) to ensure SLO compliance. This profiling enables three key capabilities: (1) It determines appropriate latency thresholds that maintain SLO compliance for various workloads and limitations (e.g., power constraints patel2024splitwise ). (2) It provides flexible adaptation by adjusting budgets based on changing workload characteristics and performance requirements. (3) It establishes a robust foundation for hybrid scheduling by accounting for both online and offline workload patterns.
4.3 SLO-aware Offline Scheduling Policies
After scheduling online requests, our offline scheduling policy repurposes the residual capacity to maximize throughput while ensuring fairness sheng2024fairness . To further optimize the serving throughput for offline requests, HyGen employs an SLO-aware Prefix Sharing Maximization (PSM) strategy. Prefix sharing is a widely adopted technique for reusing the KV cache of shared input prefixes between requests gim2024prompt ; zhao2024blendserve ; zheng2024sglang ; zheng2024batchllm .
Prefix Sharing Maximization Strategy.
Our PSM strategy organizes offline requests into a prefix tree following the structure of a Trie tree with each leaf node representing a request, capturing prefix sharing characteristics of all offline requests. The priority of each request is determined by the Depth-First Search (DFS) order of the prefix tree, where requests with the greatest prefix sharing potential are scheduled together. Subsequently, HyGen performs SLO-aware offline scheduling (Section 4.1) using this order to maximize prefix cache reuse, reducing redundant computation and improving throughput. The prefix tree structure also ensures fast insertion and deletion in runtime scheduling. A formalized algorithm is in Appendix A.2. For example, consider a system that can process two offline requests per batch with the following request queue: (What is ML, How to code, What is AI, How to debug). Under traditional FCFS scheduling, requests are processed in arrival order: (What is ML, How to code), (What is AI, How to debug), resulting in no prefix sharing opportunities. In contrast, PSM’s prefix-aware scheduling reorders requests as: (What is ML, What is AI), (How to code, How to debug), enabling KV cache reuse through shared prefixes.
The PSM strategy demonstrates strong extendability. Under certain scenarios, the vanilla PSM strategy may lead to starvation for requests with minimal prefix-sharing potential. Consider a request queue: (What is ML, What is AI, How to code, What is DL). When new requests arrive with similar prefixes (What is LLM, What is DNN), a naive prefix-sharing policy would continuously prioritize requests sharing the What is prefix, potentially starving the How to code request indefinitely. This issue can be mitigated by an extended version of our PSM policy, combining maximum prefix sharing with request freshness by using a utility ratio to ensure a balance between efficiency and fairness. Based on the utility ratio, a new offline request would be selected from the DFS order of the prefix tree or the most stale request from a self-balanced binary search tree sorted by freshness. A detailed algorithm is in Appendix A.3. These enhancements can make the PSM strategy more practical for real-world deployments, retaining its efficiency while improving fairness and adaptability.
5 Performance Evaluation
Figure 3: HyGen respects latency requirements in co-locating requests.
Figure 4: HyGen improves serving throughput under varying SLOs.
Figure 5: HyGen latency predictor achieves high accuracy for batch latency prediction.
Figure 6: Prefix Sharing Maximization improves serving throughput.
Figure 7: SLO-aware profiler contributes to HyGen’s improvements.
Figure 8: HyGen dynamically controls throughput according to online workload.
(a) SLOs
(b) Throughput
Figure 9: HyGen meets SLOs and achieves higher throughput for Yi-34B model using TP=2, PP=2.
Figure 10: HyGen meets SLO under various online QPS settings.
Figure 11: HyGen is able to meet multiple SLOs simultaneously.
(a) SLOs
(b) Throughput
Figure 12: HyGen meets SLOs and achieves higher throughput with CNN/DailyMail offline dataset.
(a) 1 hour
(b) 10 minutes
Figure 13: Request rate varies in Moonshot Mooncake’s LLM service over one-hour and ten-minute periods.
(a) SLOs
(b) Throughput
Figure 14: HyGen meets SLOs and achieves higher throughput for Mooncake trace.
(a) SLOs
(b) Throughput
Figure 15: HyGen meets SLOs and achieves higher throughput on A5000 GPU and Sheared-LLaMA-2.7B model.
Figure 16: HyGen is robust to predictor accuracy.
Figure 17: HyGen dynamically adjusts offline throughput.
5.1 Evaluation Setup
Implementation and Testbeds.
We implement HyGen on top of vLLM kwon2023efficient ; vllm2023github and Sarathi sarathi ; sarathi2024github , with 1,300 lines of additional code. We evaluate HyGen on three server configurations: one with 4 NVIDIA A100 GPUs (40GB VRAM each), one with 4 NVIDIA A40 GPUs (48GB VRAM each), and one with 1 NVIDIA A5000 GPU (24GB VRAM). All servers have 64 CPU cores, 256GB DDR4 RAM, and a 1.5TB NVMe SSD.
Models and Workloads.
For end-to-end evaluation, we use Llama2-7B touvron2023llama and Qwen-14B bai2023qwen models on A100 and A40 GPUs, respectively. Online workloads are based on the conversation trace from Azure LLM inference trace 2023 patel2024splitwise , a one-hour production trace with real-world requests and timestamps. We randomly sampled the trace to achieve the desired QPS that suits our hardware serving capacity. Specifically, within a time duration of $T$ seconds, we would sample $T\times Q$ requests to suit a desired QPS $Q$. For offline workloads, we use arXiv summarization cohan-etal-2018-discourse , a dataset for long document summarization. In our ablation studies, we evaluate HyGen across different model scales ranging from Sheared-LLaMA-2.7B xia2023sheared , Mistral-7B jiang2023mistral7b to Yi-34B young2024yi . The Mooncake trace mooncake is further used as the online trace, providing industrial request length distributions and arrival patterns, while CNN/DailyMail hermann2015teaching ; see-etal-2017-get and MMLU hendrycks2021ethics ; hendryckstest2021 are used as offline traces in the ablation studies. For interference evaluation, we focus on Time to First Token (TTFT) and Time Between Tokens (TBT), including their mean and 99th percentile (P99) values. Throughput is measured in tokens per second (TPS) and queries per second (QPS).
Baselines.
For pure online inference, we use Sarathi sarathi as our baseline. For pure offline serving, we use Sarathi-offline to evaluate the maximum offline serving capacity, where an optimal chunk size is profiled for offline workload to maximize throughput. The hyperparameter search of Sarathi-offline achieves $\sim$12% throughput gain compared to the default setup, ensuring optimal baseline performance for fair comparison. We then compare HyGen with two baselines for interference and throughput evaluation, respectively: (1) Sarathi++: We implement our online-first scheduling policy on Sarathi to support hybrid serving, including the request management and preemption handling policies introduced in Section 4.1. (2) HyGen*: To evaluate the throughput benefit of our HyGen design, we further improve Sarathi++ to an SLO-aware serving system, HyGen*. Besides inheriting the serving policies from Sarathi++, HyGen* serves offline requests at a specific offline QPS to control overall interference. The offline QPS is profiled using a similar design with the HyGen profiler to guarantee bounded SLO interference.
5.2 End-to-end Performance
HyGen respects latency requirements in co-locating requests. We evaluate HyGen under four SLO metrics (mean TBT, P99 TBT, mean TTFT, and P99 TTFT) with varied interference tolerance ratios. Figure 3 shows that HyGen controls interference and guarantees to meet specific SLOs across our settings. Compared with Sarathi++, an SLO-unaware system that yields the same result for all metrics and tolerance ratios, HyGen shows efficient SLO-aware latency control.
[…]
5.3 Performance Breakdown
Accuracy of latency predictor.
We evaluate the accuracy of our latency predictor on Llama2-7B and Qwen-14B using Azure LLM Inference trace mixed with arXiv summarization dataset. Figure 7 shows that our latency predictor achieves a mean absolute percentage error of only 1.78% and 1.07%, confirming its high accuracy.
Impact of prefix sharing.
To test HyGen’s compatibility with prefix sharing, we conducted a simulation experiment using Azure LLM Inference as the online trace and MMLU hendrycks2021ethics ; hendryckstest2021 as the offline dataset on a Llama2-7B model. In our simulation, we deducted the shared prompt prefix length between consecutive offline requests to simulate prefix sharing. Figure 7 shows that HyGen yields up to $4\times$ offline throughput gain with its prefix sharing maximization scheduling policy.
Impact of SLO-aware profiler.
To demonstrate the effect of HyGen’s SLO-aware profiler, we compared it with a simple strategy that sets the desired mean TBT SLO as the batch latency budget. Figure 7 shows the performance gap between individual batch latencies and overall mean TBT, illustrating how the SLO-aware profiler bridges this gap for controlled SLO in hybrid serving.
Breakdown by time.
Figure 9 shows a temporal throughput breakdown of HyGen. At runtime, HyGen dynamically adjusts offline throughput based on online workload and overall latency budget, batching offline requests more aggressively during online QPS troughs and reducing offline throughput during online bursts, harnessing compute resources in an adaptive manner.
5.4 Ablation Studies
Impact of model parallelisms.
To evaluate HyGen’s effectiveness in distributed inference, we deployed the Yi-34B model young2024yi on a server with 4 NVIDIA A40 GPUs using tensor-parallelism (TP) shoeybi2019megatron and pipeline-parallelism (PP) athlur2022varuna ; huang2019gpipe ; narayanan2019pipedream with degree 2 for each dimension. Using Azure LLM Inference and arXiv summarization workloads, Figure 9 shows that HyGen maintains its ability to meet SLOs and achieves higher offline throughput (up to $1.89\times$) than the baseline.
Impact of SLO requirements.
We further evaluate HyGen’s ability to meet stringent SLOs under varying online QPS settings on the four aforementioned metrics, each with 5% interference tolerance. Figure 10 shows that HyGen meets stringent SLOs for all metrics. We further demonstrate HyGen’s ability to meet multiple SLOs at the same time. By testing HyGen with a fixed P99 TTFT interference ratio (8%) and mean TBT interference ratios ranging from 10% to 50%, Figure 12 shows that at a lower mean TBT tolerance, HyGen’s performance is bounded by mean TBT SLOs; after reaching the fixed P99 TTFT SLO, mean TBT stops increasing in order to keep P99 TTFT under control.
Impact of models and datasets.
We further evaluate HyGen’s adaptability on two more experiments: The first using Mistral-7B model jiang2023mistral7b with Mooncake trace mooncake , a trace containing request lengths and timestamps taken from real-world servers, as the online trace, and arXiv summarization as the offline trace; The second experiment uses Llama2-7B model with Azure LLM Inference trace and replaced the offline dataset with CNN/DailyMail summarization dataset hermann2015teaching ; see-etal-2017-get . Figure 14 shows the varying request arrival rates of Mooncake trace over one-hour and ten-minute periods, further demonstrating the fluctuating nature of LLM services. Figure 12 and Figure 14 show that HyGen achieves superior performance than its counterparts under these settings.
Impact of hardware testbeds.
[…]
Impact of predictor accuracy.
[…]
Impact of online arrival rate.
Figure 17 shows the effect of online QPS on offline throughput with 5% P99 TBT tolerance. As online load increases, HyGen adjusts offline throughput based on the system’s residual serving capacity while maintaining higher throughputs. Understandably, a high online arrival rate limits the headroom for co-location as it approaches system serving capacity.
6 Conclusion
[…]
Acknowledgements
We thank the anonymous reviewers for their constructive and insightful feedback. This work was supported in part by grants from Cisco and Google, and by an award from NVIDIA Academic Program. It also utilized the Delta system at the National Center for Supercomputing Applications (NCSA) through allocation CIS240236 from the ACCESS program.
Appendix
Appendix A Algorithms
A.1 HyGen Two-phase Scheduling Algorithm
This section gives a detailed and formalized demonstration of the two-phase scheduling algorithm introduced in Section 4.1. In each scheduling step, the workflow invokes the SLO-aware scheduling process in Algorithm 1 twice (line 13 and line 18) to form a hybrid batch with online and offline request co-location while respecting latency and memory limits. To reduce the scheduling overhead of HyGen, we employ a message queue for asynchronous communication between the main process and the offline scheduler. After each scheduling step, the main process sends the metadata of batched requests to the message queue. The offline scheduler first calculates the expected status of each request based on scheduling decisions from the previous batch, and then runs a scheduling simulation to generate offline request scheduling decisions using our latency predictor and the profiled latency budget. The offline scheduling decisions are then sent back to the main process using the message queue and used for the next scheduling step. To support pipeline parallelization, a scheduling history archive of $K$ steps is kept by the offline scheduler for pipeline parallelization degree $K$, in order to have a holistic view of every request running in each pipeline stage at the time.
Algorithm 2 HyGen two-phase scheduler
1:global $Q_{on}$ (online request queue), $Q_{off}$ (offline request queue)
2:global $R_{on}$ (online request running list), $R_{off}$ (offline request running list)
3:global $Q_{send}$, $Q_{recv}$ (message queues)
4:global $L$ (latency budget), $M$ (memory budget), $C$ (chunk size), $M_{off}$ (offline memory)
5:function ASYNC_SCHEDULER
6:  while True do
7:   scheduled requests $S\leftarrow Q_{send}$.get(block=True)
8:   UPDATE_REQUEST_STATUS($S$, $R_{on}$, $R_{off}$)
9:   batched requests $B\leftarrow$ {}
10:   latency budget $t\leftarrow L$
11:   memory budget $m\leftarrow$ GET_FREE_MEMORY() + $M_{off}$
12:   chunk size $c\leftarrow C$
13:   $B\leftarrow B\cup$ SLO_AWARE_SCHEDULE($R_{on}$, $Q_{on}$, $t$, $c$, $m$)
14:   if $m<M_{off}$ then
15:    PREEMPT_OFFLINE($R_{off}$, $m$, $M_{off}$)
16:   $m\leftarrow m-M_{off}$
17:   $B\leftarrow B\cup$ SLO_AWARE_SCHEDULE($R_{off}$, $Q_{off}$, $t$, $c$, $m$)
18:   $Q_{recv}$.send($B$, $t$, $m$)
19:function SCHEDULER
20:  Output: batched requests $B$
21:  scheduled requests $B$, latency budget $t$, memory budget $m$ $\leftarrow$ $Q_{recv}$.get(block=True)
22:  if REQUEST_ARRIVAL then
23:   UPDATE_BUDGET($B$, $t$, $m$)
24:   if $t<0$ or $m<0$ then
25:    PREEMPT_UNTIL_FIT($B$, $t$, $m$)
26:  $Q_{send}$.send($B$)
27:  return $B$
A.2 SLO-Aware Prefix Sharing Maximization Algorithm
This section details the SLO-aware prefix sharing maximization algorithm design in HyGen. We construct a prefix tree $T_{p}$ to capture prefix sharing characteristics among offline requests. During scheduling, offline requests are selected in the DFS order of the prefix tree, and deleted once being scheduled. Additionally, running requests keep their original DFS order in future scheduling process, effectively utilizing prefix sharing.
Algorithm 3 Prefix-sharing-aware offline scheduler
1: Construct prefix tree $T_{p}$
2:function PREFIX_SHARING_OFFLINE_SCHEDULE
3:  Input: running requests $R$, latency budget $t$, remaining chunk size $c$, memory budget $m$
4:  Output: batched requests $B$
5:  $B\leftarrow\{\}$
6:  for $r\in R$ do
7:   if $r$.state == DECODE then
8:    $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
9:    if $t>t_{req}$ then
10:     break
11:    $t\leftarrow t-t_{req}$
12:    $B\leftarrow B\cup$ {($r$, 0, $t_{req}$)}
13:   else if $r$.state == PREFILL then
14:    $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
15:    if $l>0$ then
16:     $t\leftarrow t-t_{req}$
17:     $c\leftarrow c-l$
18:     $m\leftarrow m$- GET_NUM_BLOCKS($l$)
19:     $B\leftarrow B\cup$ {($r$, $l$, $t_{req}$)}
20:    else
21:     break
22:  while $T_{p}$ is not empty do
23:   $r\leftarrow T_{p}$.get_next_request()
24:   $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
25:   if $l>0$ then
26:    $t\leftarrow t-t_{req}$
27:    $c\leftarrow c-l$
28:    $m\leftarrow m$- GET_NUM_BLOCKS($l$)
29:    $B\leftarrow B\cup$ {($r$, $l$, $t_{req}$)}
30:    $T_{p}$.remove_request($r$)
31:   else
32:    break
33:  return $B$
A.3 Extended SLO-Aware Prefix Sharing Maximization Algorithm
This section details the extended version of the SLO-aware prefix sharing maximization algorithm, enhanced with fairness-aware scheduling. For offline requests, we construct a prefix tree $T_{p}$ for prefix sharing, and a self-balancing BST $T_{f}$ for request freshness. A utility ratio between 0 and 1 is used to balance the chance between these two data structures. During scheduling, the extended prefix sharing maximization algorithm retrieves waiting offline requests from either $T_{p}$ or $T_{f}$, based on the utility ratio. The selected request is then deleted from both data structures to ensure synchronization. This solution balances prefix sharing and fairness without disrupting the DFS order of the prefix tree while avoiding possible starvation.
Algorithm 4 Prefix-sharing-aware offline scheduler
1: Construct prefix tree $T_{p}$ and self-balanced BST $T_{f}$
2:function PREFIX_SHARING_OFFLINE_SCHEDULE
3:  Input: running requests $R$, latency budget $t$, remaining chunk size $c$, memory budget $m$, utility value $u$
4:  Output: batched requests $B$
5:  $B\leftarrow\{\}$
6:  for $r\in R$ do
7:   if $r$.state == DECODE then
8:    $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
9:    if $t>t_{req}$ then
10:     break
11:    $t\leftarrow t-t_{req}$, $B\leftarrow B\cup$ {($r$, 0, $t_{req}$)}
12:   else if $r$.state == PREFILL then
13:    $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
14:    if $l>0$ then
15:     $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m$- GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$ {($r$, $l$, $t_{req}$)}
16:    else
17:     break
18:  while $T_{p}$ is not empty do
19:   $rand\leftarrow$ RANDOM_NUMBER(0, 1)
20:   if $rand<u$ then
21:    $r\leftarrow T_{p}$.get_next_request()
22:   else
23:    $r\leftarrow T_{f}$.get_next_request()
24:   $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
25:   if $l>0$ then
26:    $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m$- GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$ {($r$, $l$, $t_{req}$)}
27:    $T_{p}$.remove_request($r$)
28:    $T_{f}$.remove_request($r$)
29:   else
30:    break
31:  return $B$
A.4 Complexity Analysis of the Two-Phase Scheduling Algorithm
In this section, we present an analysis of the computational complexity of HyGen’s two-phase scheduler. The time complexity of its core components is as follows:
  • Latency Prediction: $O(1)$ inference using a pre-trained LR model.
  • PSM: The initial construction of the prefix tree is $O(N×L)$, where $N$ is the number of offline requests and $L$ is the average number of tokens. Each insertion or deletion costs $O(L)$. In implementation, getting the next request in the DFS order only takes $O(1)$, since the DFS order is put in a pre-processed list derived from the prefix tree and can be synced up with the prefix tree asynchronously.
  • PSM with Fairness: In the fairness-aware PSM algorithm, a self-balancing BST is used for picking the stalest request. Each lookup, insertion, or deletion takes $O(\log n)$. In implementation, the requests can be kept in an FCFS queue, which syncs up with the BST asynchronously to guarantee correctness, so that each lookup still only takes $O(1)$ time.
Overall, for the two-phase scheduling using existing policies, the time complexity remains $O(n)$ as asynchronous updates can be performed for advanced policies like fairness-aware PSM.
Appendix B Further Discussion of the Latency Predictor
Effective co-location in HyGen hinges on a latency predictor that is both highly accurate and computationally lightweight, as it enables real-time scheduling decisions without introducing significant overhead. To this end, we employ a linear regression (LR) model, which provides inference in constant time, $O(1)$, making it ideal for a dynamic serving environment. This section details the model’s formulation and its adaptability to the complexities of real-world deployment scenarios.
Model Formulation.
The execution time of a serving batch, $T_{\text{batch}}$, is primarily determined by the computational patterns of its prefill and decode stages. We model this relationship as a function of the batch’s composition:
$T_{\text{batch}}=f(S_{p},S_{d},S_{p}^{2},N_{p},N_{d})$ (2)
where $S_{p}$ and $S_{d}$ represent the total number of tokens in the prefill and decode phases, respectively, and $N_{p}$ and $N_{d}$ are the corresponding request counts. The quadratic term $S_{p}^{2}$ is crucial for capturing the non-linear scaling of the self-attention mechanism, which dominates the computational cost of the prefill stage. In contrast, the decode stage, which processes one token per request at a time, exhibits linear scaling with the batch size ($N_{d}$). The model’s coefficients are pre-trained on data gathered by systematically profiling the target hardware and LLM across a diverse set of batch compositions, a lightweight process that ensures applicability to any deployment environment.
Robustness Through System-Level Design.
While a linear model offers unparalleled efficiency, its accuracy in the face of dynamic system conditions (e.g., GPU temperature variations, resource contention) is a critical consideration. HyGen ensures robustness not through a more complex model, but through a synergistic system design. The SLO-aware profiler (Section 4.2) first establishes a macro-level operational latency budget, which implicitly captures the system’s current performance characteristics. The LR predictor then operates at a micro-level, making fine-grained decisions on batch composition within this pre-calibrated budget.
This two-level approach effectively decouples the real-time scheduling decision from low-level hardware variability. As shown in Figure 17, this design allows the system to maintain robust performance and meet SLOs even in scenarios with predictor error rates exceeding 20%. Furthermore, the feature set is extensible; environmental factors like hardware load can be readily integrated into the model if required for specific use cases.
Adaptability to Modern Model Architectures.
The feature set defined in Equation 2 is sufficiently general to generalize across various modern LLM architectures without modification. For instance, for Mixture-of-Experts (MoE) models [48], where a fixed number of experts are activated per token, the resulting computational cost scales linearly with the total number of tokens processed and is effectively modeled by the $S_{p}$ and $S_{d}$ features. Similarly, in hybrid architectures that combine linear-complexity components (e.g., Mamba [19]) with quadratic-complexity attention, our model naturally captures both computational patterns; the linear cost is reflected in the learned coefficient for $S_{p}$, while the quadratic cost of the Transformer blocks is captured by the $S_{p}^{2}$ term. This adaptability underscores the predictor’s robust design, ensuring its relevance as LLM architectures continue to evolve.
Appendix C Taming the Throughput-Latency Tradeoff in LLM Serving with HyGen
Traditional instance scaling solutions address bursty workloads by launching new instances, which can cause tens of seconds to several minutes cold-start delays [15, 64, 66]. To handle these delays, providers often keep standby instances online, leading to wasted resources during off-peak periods.
In contrast, HyGen optimizes resource utilization by running offline workloads on idle resources and reallocating them to online requests in real-time, within a single inference iteration. This eliminates cold-start delays while ensuring high resource utilization.
HyGen complements instance scaling solutions by automating the transition between online and offline workloads, reducing the need for manual intervention. While instance scaling manages large load fluctuations, HyGen ensures efficient resource use during low-traffic periods, optimizing overall system performance in fixed-size clusters.